### v1

# Adversarially Perturbed Precision Matrix Estimation

Abstract
Precision matrix estimation is a fundamental topic in multivariate statistics and modern machine learning. This paper proposes an adversarially perturbed precision matrix estimation framework, motivated by recent developments in adversarial training. The proposed framework is versatile for the precision matrix problem since, by adapting to different perturbation geometries, the proposed framework can not only recover the existing distributionally robust method but also inspire a novel moment-adaptive approach to precision matrix estimation, proven capable of sparsity recovery and adversarial robustness. Notably, the proposed perturbed precision matrix framework is proven to be asymptotically equivalent to regularized precision matrix estimation, and the asymptotic normality can be established accordingly. The resulting asymptotic distribution highlights the asymptotic bias introduced by perturbation and identifies conditions under which the perturbed estimation can be unbiased in the asymptotic sense. Numerical experiments on both synthetic and real data demonstrate the desirable performance of the proposed adversarially perturbed approach in practice.
1 Introduction
The precision matrix, also known as the inverse covariance matrix, can reveal the conditional correlations between pairs of variables given all other variables (Fan et al., 2016). Estimating the precision matrix is essential in many problems, including neuroimaging (Huang et al., 2010), network analysis (Torri et al., 2018), portfolio optimization (Procacci and Aste, 2022), and genomics (Schäfer and Strimmer, 2005), among others.
Suppose $\bm{x}_{1},\ldots,\bm{x}_{n}$ are $n$ independent and identically distributed samples drawn from a $d$-dimensional multivariate normal distribution. For simplicity, we assume that the population mean is zero and the population covariance matrix $\Sigma$ is positive definite and unknown. The standard approach to obtain the precision matrix estimation, i.e., $C=\Sigma^{-1}$, is to identify the minimizer of the negative log-likelihood function:
$\min_{C\succ 0}-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}[\bm{x}^{\top}C\bm{x}],$ (1)
where $C\succ 0$ means that $C$ is positive definite, $\mathrm{det}C$ denotes the determinant of the matrix $C$, $\mathbb{P}_{n}$ is the empirical distribution. When the matrix dimension is small such that $d<n$, the solution to the maximum likelihood estimation problem is the inverse of the sample covariance matrix $\mathbb{E}_{\mathbb{P}_{n}}[\bm{x}\bm{x}^{\top}]$. However, if the matrix dimension is large such that $d>n$, the maximum likelihood estimation problem does not admit a solution. Numerous studies have been conducted to explore shrinkage and regularization frameworks for precision matrix estimation to achieve estimations which are invertible, well-conditioned, and enjoy desirable statistical properties such as sparsity (Yuan and Lin, 2007; Ledoit and Wolf, 2012; Nguyen et al., 2022). We will show later that the proposed adversarially perturbed framework in this paper can recover some of these shrinkage and regularized effects by selecting different types of perturbations.
Our proposed adversarially perturbed precision matrix framework is inspired by recent developments in trustworthy machine-learning—adversarial training (Goodfellow et al., 2015). The perturbed estimation is obtained by solving the following optimization problem:
$\inf_{C\succ 0}\left\{-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}\left[\max_{\|\bm{\Delta}\|\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C(\bm{x}+\bm{\Delta})\right]\right\},$ (2)
where $\|\cdot\|$ is the norm of the perturbation $\bm{\Delta}$, and $\delta$ is the magnitude of the perturbation. Instead of minimizing the empirical average of the likelihood function, the adversarially perturbed estimation problem (2) seeks for the minimizer of the loss under the most adversarial perturbation $\bm{\Delta}$, which is determined by both $C$ and $\bm{x}$ through the inner maximization problem in (2), subject to the constraint that the norm of the perturbation is bounded.
In this paper, we will focus on the $\ell_{p}$-norm of the perturbation. When $p=2$, it can be proven that the perturbed estimation problem (2) is equivalent to the Wasserstein shrinkage estimation proposed in Nguyen et al. (2022). The resulting optimization problem is tractable and yields a nonlinear shrinkage estimator that is well-conditioned, rotation equivalent, and preserves the eigenvalue order (Nguyen et al., 2022). When $p=\infty$, by optimizing a convex upper-bound surrogate of the objective function in the outer minimization in problem (2), the resulting estimation can be proven to promote sparsity and adversarial robustness.
We demonstrate that introducing perturbations is asymptotically equivalent to imposing regularization in precision matrix estimation, provided the perturbation magnitude is small. The corresponding first-order and second-order regularization terms are also given. Based on the regularization effect, the asymptotic distribution of the adversarially perturbed estimation is derived. The asymptotic distribution characterizes the bias induced by the inner maximization problem in our proposed framework (2). We suppose the perturbation magnitude follows the rule $\delta=\eta/n^{\gamma}$, where $\eta>0$ and $\gamma>0$. Our analysis yields the following two regimes: When $\gamma>1/2$, the asymptotic bias is zero. In this regime, the adversarial perturbation decays sufficiently fast such that it does not affect estimation accuracy. When $\gamma\leq 1/2$, the asymptotic distribution has a non-zero bias. Furthermore, specifically for $\ell_{\infty}$-norm perturbations using the upper-bound surrogate of the objective function, the derived asymptotic distribution demonstrates that sparsity recovery is achievable when $\gamma=1/2$. The regularization reformulation and asymptotic distribution provide insights into the statistical behavior of both Wasserstein shrinkage estimation and the sparse estimation recovered by our novel framework. See a summary table of the associated asymptotic behavior in Table 1, Section 5.
Numerical experiments are delivered for the case when $p=\infty$ and the surrogate objective is considered. We carry out the numerical experiments on both synthetic data and a real-world dataset. Firstly, the numerical experiments on the synthetic data validate our theoretical results that the perturbed estimation can recover the sparsity and achieve adversarial robustness. Secondly, by adopting linear discriminant analysis based on precision matrix estimation for the real dataset, our method shows more desirable practical performance.
1.1 Related Work
1.1.1 Adversarial Training
As the modern artificial intelligence and machine learning models grow more powerful and widely used, their reliability becomes increasingly important. One major challenge to reliability is that machine learning models can be vulnerable to small changes in input data, which may lead to unreliable outputs (Goodfellow et al., 2015). To improve the robustness of the models, adversarial training has been introduced as a principled approach by optimizing against adversarial perturbations posed in the input data (Goodfellow et al., 2015; Madry et al., 2018).
Different from the traditional empirical risk minimization approach, the resulting adversarial training problem is a min-max robust optimization problem and thus introduces a new statistical estimation paradigm. While most of the relevant existing work focus on the vector-valued parameter (Goodfellow et al., 2015; Madry et al., 2018; Javanmard and Soltanolkotabi, 2022; Ribeiro et al., 2023; Xie and Huo, 2024, 2025), our proposed framework provides novel insights on how the adversarial training motivates a new estimation framework for the matrix-valued parameter. The proposed framework serves as an alternative for practitioners when the precision matrix estimation with desirable properties such as sparsity and adversarial robustness is needed.
Adversarial training is closely related to the distributionally robust optimization. It is well known that adversarial training can be viewed as a special case of the Wasserstein distributionally robust optimization (Staib and Jegelka, 2017; Gao et al., 2024). This paper, however, provides another perspective by demonstrating that, in the context of precision matrix estimation, the Wasserstein distributionally robust optimization is a special case of adversarial training.
1.1.2 Precision Matrix Estimation
Precision matrix estimation has been extensively studied under different settings. Classical approaches primarily focus on enforcing sparsity via $\ell_{1}$-regularization or $\ell_{1}$-minimization (Yuan and Lin, 2007; Friedman et al., 2008; Cai et al., 2011). As applications have broadened, researchers have developed methods robust to data anomalies. For instance, Loh and Tan (2018) establishes a framework robust to data contamination, Tang et al. (2021) focuses on mitigating the impact of outliers, Avella-Medina et al. (2018) studies robustness against heavy-tailed distributions, and the method framed in Nguyen et al. (2022) addresses distributional shifts.
Our framework can recover the Wasserstein shrinkage estimation (Nguyen et al., 2022) under $\ell_{2}$-perturbation while being capable of the sparsity recovery achieved by classical $\ell_{1}$-regularized methods (Yuan and Lin, 2007; Friedman et al., 2008). In addition, by defining the model through an adversarial training lens, we achieve intrinsic adversarial robustness. Specifically, as detailed in Section 6.1, under $\ell_{\infty}$-perturbation and a specific surrogate objective, our formulation reduces to a moment-adaptive weighted $\ell_{1}$-regularized precision matrix estimation problem, thereby simultaneously achieving sparse recovery and adversarial robustness.
1.2 Notations and Definitions
$\bm{x}]_{i}$ denotes the $i$th component of vector $\bm{x}$. $\operatorname{sign}([\bm{x}]_{i})$ denotes the sign of $\bm{x}]_{i}$ (where we let $\operatorname{sign}(0)=0$), and accordingly we let $\operatorname{sign}(\bm{x})=(\operatorname{sign}([\bm{x}]_{1}),...,\operatorname{sign}([\bm{x}]_{d}))^{\top}$. We use $\|\bm{x}\|_{p}$ to denote the $\ell_{p}$-norm of the vector $\bm{z}$, i.e., $\|\bm{x}\|_{p}^{p}=\sum_{j=1}^{d}|x_{j}|^{p},1\leq p<\infty,\|\bm{x}\|_{\infty}=\max_{1\leq j\leq d}|x_{j}|.$ $\odot$ denotes the component-wise product. $\mathbb{I}(\cdot)$ denotes the indicator function. $\mathrm{vec}(X)$ denotes the vectorization of the matrix $X$.
1.3 Organization of this Paper
The remainder of this paper is organized as follows. In Section 2, we study the adversarially $\ell_{2}$-perturbed precision matrix estimation and its equivalence to the Wasserstein shrinkage estimation. In Section 3, we investigate the formulation and the properties of the adversarially $\ell_{\infty}$-perturbed precision matrix estimation. In Section 4, we introduce the regularization effect of the proposed framework. In Section 5, we study the asymptotic distribution of the proposed precision matrix estimation. Numerical experiments are conducted and analyzed in Section 6. Future work is discussed in Section 7. The proofs are relegated to the Supplementary Material whenever possible.
2 Adversarially $\ell_{2}$-perturbed Precision Matrix Estimation
In this section, we will discuss the $\ell_{2}$-perturbed precision matrix estimation.
The adversarially perturbed precision matrix estimation under $\ell_{2}$-norm is the solution to the following optimization problem:
$\inf_{C\succ 0}\left\{-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}\left[\max_{\|\bm{\Delta}\|_{2}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C(\bm{x}+\bm{\Delta})\right]\right\},$ (3)
where the perturbation norm in the constraint set in the inner maximization problem is chosen as $\ell_{2}$ norm.
2.1 Tractable Reformulation
In this subsection, we give the convex tractable reformulations of the adversarially $\ell_{2}$-perturbed precision matrix estimation.
Theorem 2.1 (Convex Reformulations under $\ell_{2}$-perturbation).
The problem (3) is equivalent to the following problem:
$\inf_{C\succ 0,\lambda I-C\succ 0}\left\{-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}[\bm{x}^{\top}C\bm{x}]+\lambda\delta^{2}+\mathbb{E}_{\mathbb{P}_{n}}[\bm{x}^{\top}C(\lambda I-C)^{-1}C\bm{x}]\right\},$ (4)
which further admits the following reformulation:
$\inf_{\lambda,\{t_{i}\},C}$ $-\log\mathrm{det}C+\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}^{\top}C\bm{x}_{i}+\lambda\delta^{2}+\frac{1}{n}\sum_{i=1}^{n}t_{i}$ (5)
$\mathrm{subject\ to}$ $C\succ 0,\lambda I-C\succ 0,$
$\begin{pmatrix}t_{i}&\bm{x}_{i}^{\top}C\\ C\bm{x}_{i}&\lambda I-C\end{pmatrix}\succeq 0,\quad\forall i=1,\dots,n.$
The reformulation (4) is obtained by deriving the dual of the inner maximization problem in (3), which admits a closed-form solution under $\ell_{2}$-perturbation. Subsequently, applying the Schur complement yields formulation (5).
Problem (5) is a convex semidefinite optimization problem with a log-determinant barrier term in the objective function and linear matrix inequalities in the decision variables $\lambda,\{t_{i}\},C$. Such problems fall within the framework of semidefinite (conic) programming and can be solved efficiently by interior-point methods (Boyd and Vandenberghe, 2004; Nemirovski, 2007).
2.2 Equivalence to the Wasserstein Shrinkage Estimation
In this subsection, we will show that the $\ell_{2}$-perturbed precision matrix estimation problem is equivalent to the Wasserstein distributionally robust precision matrix estimation proposed in Nguyen et al. (2022), which they coin as the Wasserstein shrinkage estimation.
The optimization problem for obtaining the Wasserstein shrinkage estimation introduced in Nguyen et al. (2022) can be written as the following problem:
$\inf_{C\succ 0}\left\{-\log\mathrm{det}C+\sup_{S\in\mathcal{U}_{\bar{A},\rho}}\mathrm{tr}(SC),\right\},$ (6)
where $\mathcal{U}_{\bar{A},\rho}$ is the uncertainty set and $\rho$ is the radius of the set. The uncertainty set is centered at the sample covariance $\bar{A}=\mathbb{E}_{\mathbb{P}_{n}}[\bm{x}\bm{x}^{\top}]$ and is defined in terms of the metric induced by the 2-Wasserstein distance on the family of normal distributions, and we have that
$\mathcal{U}_{\bar{A},\rho}=\left\{S\succeq 0:\mathrm{tr}(\bar{A})+\mathrm{tr}(S)-2\mathrm{tr}(\sqrt{\bar{A}^{\frac{1}{2}}S\bar{A}^{\frac{1}{2}}})\leq\rho^{2}\right\}.$
Nguyen et al. (2022) shows that the min-max optimization problem in (6) admits a convex reformulation, as restated in Theorem 2.2.
Theorem 2.2 (Nguyen et al. (2022)).
For any fixed $\rho>0$, the Wasserstein distributionally robust estimation problem (6) is equivalent to
$\inf_{C\succ 0,\lambda I-C\succ 0}\left\{-\log\mathrm{det}C+\lambda\left(\rho^{2}-\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}\bm{x}_{i}^{\top}\right)+\lambda^{2}\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}(\lambda I-C)^{-1}\bm{x}_{i}^{\top}\right\}.$ (7)
The equivalence of our adversarially perturbed estimation problem (3) and distributionally robust trained problem (6) can be built upon problem (4) and (7) based on algebraic calculations. We state this result in the following theorem.
Theorem 2.3 (Equivalence).
For fixed $\rho=\delta$, the $\ell_{2}$-perturbed precision matrix estimation problem (3) is equivalent to the Wasserstein distributionally robust precision matrix estimation problem (6).
Theorem 2.3 implies that, in the precision matrix estimation problem, our proposed adversarially perturbed approach could recover the Wasserstein distributionally estimator if the $\ell_{2}$-perturbation is chosen. Imposing an $\ell_{2}$-perturbation to the observation $\bm{x}$ is equivalent to constructing an ambiguity set for the sample covariance matrix in terms of the Wasserstein distance. As illustrated in Nguyen et al. (2022), the resulting estimator is invertible, rotation equivalent, preserves the eigenvalue order, and can be obtained by solving a sequential quadratic approximation algorithm.
Although arriving at the same solution, our proposed perturbed framework (3) adopts a sample-wise perturbation perspective. Crucially, this formulation admits a tractable asymptotic expansion, from which the asymptotic distribution is derived, seeing Section 5. This framework elucidates the statistical effect of shrinkage and enables statistical inference.
3 Adversarially $\ell_{\infty}$-perturbed Precision Matrix Estimation
In this section, we explore the adversarially $\ell_{\infty}$-perturbed precision matrix estimation.
For the adversarially perturbed precision matrix estimation (2), if the norm of the perturbation is chosen as $\ell_{\infty}$-norm, we have the following formulation:
$\inf_{C\succ 0}\left\{-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}\left[\max_{\|\bm{\Delta}\|_{\infty}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C(\bm{x}+\bm{\Delta})\right]\right\}.$ (8)
3.1 Tractable Surrogate
Notably, the inner maximization problem in problem (8) is NP-hard (Nemirovski et al., 1999). To ensure computational tractability, we replace the NP-hard inner maximization in (8) with a convex upper-bound surrogate. Based on the triangle inequality on the perturbation terms, we derive the following formulation:
$\inf_{C\succ 0}\left\{-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{x}^{\top}C\bm{x}\right]+2\delta\sum_{j=1}^{d}\sum_{k=1}^{d}\omega_{k}|[C]_{kj}|+\delta^{2}\|C\|_{1,1}\right\}$ (9)
where
$\omega_{k}=\mathbb{E}_{\mathbb{P}_{n}}\left[\left|[\bm{x}]_{k}\right|\right].$
The formulation (9) can be interpreted as a regularized precision matrix estimation with penalties with orders $O(\delta)$ and $O(\delta^{2})$. The term $2\delta\sum_{j,k}\omega_{k}|[C]_{kj}|$ scales with $O(\delta)$ and functions as a moment-adaptive weighted $\ell_{1}$-regularizer. The weights $\omega_{k}$’s enforce a heterogeneous shrinkage scheme where entries associated with variables of larger average magnitude are penalized more heavily. Since high-magnitude features act as high-leverage points for adversarial exploitation, this weighting scheme promotes an estimation that is not only sparse but also intrinsically robust to adversarial manipulation. The quadratic term $\delta^{2}\|C\|_{1,1}$ addresses the second-order interactions of the adversarial perturbation. By penalizing the entry-wise magnitude of matrix $C$, this quadratic regularizer serves as a global stability adjustment, ensuring the surrogate remains a rigorous upper bound for the original objective in (8).
3.1.1 Properties
We will discuss the properties of the estimation obtained from (9), including the adversarial robustness stated in Proposition 3.1 and the sparsity recovery stated in Corollary 3.2.
Proposition 3.1 (Adversarial Robustness).
Suppose $\widetilde{C}$ is the minimizer of problem (9), and $\mathcal{S}(\cdot)$ is the associated objective function. Then, the adversarial robust loss of $\widetilde{C}$ is upper bounded, i.e.,
$\log\mathrm{det}\widetilde{C}+\mathbb{E}_{\mathbb{P}_{n}}\left[\max_{\|\bm{\Delta}\|_{p}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}\widetilde{C}(\bm{x}+\bm{\Delta})\right]\leq\inf_{C\succ 0}\mathcal{S}(C),\quad 1\leq p\leq\infty.$
Proposition 3.1 implies that the adversarially robust loss of the estimation $\widetilde{C}$ obtained from problem (9) under $\ell_{p}$-perturbation, which is defined as the worst-case loss under adversarial perturbations posed to the samples, is upper bounded. This property is due to the nature of adversarial training, the motivation behind our proposed problem, and the way we formulate problem (9) as an upper-bound surrogate.
Corollary 3.2 (Sparsity Recovery).
Let $\widetilde{C}$ be the minimizer of problem (9). Suppose that $\delta=\eta/\sqrt{n}$ with $\eta>0$. Then, as $n\to\infty$, the asymptotic distribution of $\widetilde{C}]_{ij}$ has a positive probability mass at 0 whenever $\Sigma^{-1}]_{ij}=0$.
Remark 3.3.
Corollary 3.2 follows immediately from Theorem 5.2 in Section 5 and Proposition 3.8 in Xie and Huo (2025).
Corollary 3.2 implies that for zero entries in the true precision matrix $\Sigma^{-1}$, the asymptotic distribution of the corresponding estimates from (9) has a positive probability at zero. We call this property sparsity recovery, which is theoretically guaranteed if the adversarial perturbation order is $1/\sqrt{n}$. The sparsity is induced by the first penalty term in (9). Notably, while similar to the shrinkage achieved in $\ell_{1}$-regularized precision matrix estimation (Yuan and Lin, 2007; Friedman et al., 2008), this term incorporates the adaptive adjustment for adversarial robustness discussed earlier
4 Regularization Effect
We discuss the regularization effect induced by the adversarial perturbation in the proposed framework.
We characterize this regularization effect mathematically in the following proposition.
Proposition 4.1 (Regularization Effect).
Suppose the random variable $\bm{x}$ follows a distribution $\bm{P}$ with $\mathbb{E}_{\bm{P}}[\|\bm{x}\|_{p}]<\infty$ and $\bm{P}([C\bm{x}]_{i}=0)=0$, $C$ is a positive definite matrix with $\|C\|_{p\to q}<\infty$, and $p\in(1,\infty]$ with $1/p+1/q=1$. Then, we have that
$\mathbb{E}_{\bm{P}}\left[\max_{\|\bm{\Delta}\|_{p}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C(\bm{x}+\bm{\Delta})\right]=\mathbb{E}_{\bm{P}}[\bm{x}^{\top}C\bm{x}]+2\delta\mathbb{E}_{\bm{P}}\left[\|C\bm{x}\|_{q}\right]+\delta^{2}\mathbb{E}_{\bm{P}}\left[\bm{v}_{C\bm{x}}^{\top}C\bm{v}_{C\bm{x}}\right]+o(\delta^{2}),$
as $\delta\to 0$, where
$\bm{v}_{C\bm{x}}=\frac{\operatorname{sign}(C\bm{x})\odot|C\bm{x}|^{q-1}}{\|C\bm{x}\|_{q}^{q-1}}.$ (10)
Remark 4.2.
Since the term $\log\det C$ in the proposed framework (2) is deterministic, Proposition 4.1 focuses on the inner maximization problem in (2) to analyze the induced regularization under $\ell_{p}$-perturbation. Instead of only analyzing on the empirical distribution, we consider a more general distribution $\bm{P}$ satisfying certain regularity conditions. The stated regularity conditions in Proposition 4.1 ensure well-posedness and integrability.
Example 4.3.
We provide the examples for the cases when $p=2$ and $p=\infty$ in Proposition 4.1:
$\mathbb{E}_{\bm{P}}\left[\max_{\|\bm{\Delta}\|_{\infty}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C(\bm{x}+\bm{\Delta})\right]$
$=$ $\mathbb{E}_{\bm{P}}[\bm{x}^{\top}C\bm{x}]+2\delta\mathbb{E}_{\bm{P}}\left[\|C\bm{x}\|_{1}\right]+\delta^{2}\mathbb{E}_{\bm{P}}\left[\operatorname{sign}(C\bm{x})^{\top}C\operatorname{sign}(C\bm{x})\right]+o(\delta^{2}),$
$\mathbb{E}_{\bm{P}}\left[\max_{\|\bm{\Delta}\|_{2}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C(\bm{x}+\bm{\Delta})\right]$
$=$ $\mathbb{E}_{\bm{P}}[\bm{x}^{\top}C\bm{x}]+2\delta\mathbb{E}_{\bm{P}}\left[\|C\bm{x}\|_{2}\right]+\delta^{2}\mathbb{E}_{\bm{P}}\left[\frac{\bm{x}C^{3}\bm{x}}{\bm{x}C^{2}\bm{x}}\right]+o(\delta^{2}),$
as $\delta\to 0$.
Proposition 4.1 reveals that the proposed adversarially robust precision matrix framework is equivalent to the regularized precision matrix estimation in the asymptotic sense. For small perturbations, the worst-case loss inflates by a first-order term $2\delta\mathbb{E}_{\bm{P}}[\|C\bm{x}\|_{q}]$, a second-order term $\delta^{2}\mathbb{E}_{\bm{P}}[\bm{v}_{C\bm{x}}^{\top}C\bm{v}_{C\bm{x}}]$, and a higher-order residual.
The first-order regularization term arises from Hölder’s inequality and penalizes the $\ell_{q}$ norm of the gradient direction, where $1/p+1/q=1$. The vector $\bm{v}_{C\bm{x}}$ in the second-order term is the $\ell_{p}$-unit vector that achieves the associated Hölder dual pairing, as defined in (10). The second-order regularization term reflects the local curvature of the worst-case loss. Specifically, this term corresponds to the mean of the Rayleigh quotient of $C$ evaluated with respect to the vector $C\bm{x}$ (when $p=2$) or $\operatorname{sign}(C\bm{x})$ (when $p=\infty$). If $C$ has the eigen-decomposition $C=\sum_{j}\lambda_{j}\bm{u}_{j}\bm{u}_{j}^{\top}$, the associated Rayleigh quotient represents a weighted average of the eigenvalues. For $p=2$, it takes the form:
$\frac{\bm{x}^{\top}C^{3}\bm{x}}{\bm{x}^{\top}C^{2}\bm{x}}=\sum_{j}\lambda_{j}\left(\frac{\lambda_{j}^{2}(\bm{u}_{j}^{\top}\bm{x})^{2}}{\sum_{k}\lambda_{k}^{2}(\bm{u}_{k}^{\top}\bm{x})^{2}}\right).$ (11)
Similarly, for $p=\infty$:
$\operatorname{sign}(C\bm{x})^{\top}C\operatorname{sign}(C\bm{x})=\sum_{j}\lambda_{j}(\bm{u}_{j}^{\top}\operatorname{sign}(C\bm{x}))^{2}.$
Thus, the second-order regularization implies that adversarial perturbation induces a shrinkage effect on the eigenvalues of $C$. Furthermore, as seen in (11) for the $p=2$ case, this shrinkage effect is stronger on larger eigenvalues due to the $\lambda^{2}$ weighting.
We could establish the asymptotic distribution of the adversarially perturbed precision estimation utilizing the regularization effect in Proposition 4.1.
5 Asymptotic Distribution
This section will provide the asymptotic distribution of the adversarially perturbed estimator under different settings.
Recall that the proposed adversarially perturbed estimation is the solution to the following optimization problem:
$\widehat{C}\in\arg\inf_{C\succ 0}\left\{-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}\left[\max_{\|\bm{\Delta}\|_{p}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C(\bm{x}+\bm{\Delta})\right]\right\}.$ (12)
Regarding the perturbation magnitude, we consider a regime where the perturbation level $\delta$ is small and decays as the sample size $n$ increases. Formally, we let $\delta=\delta_{n}=\eta n^{-\gamma}$ with constants $\eta,\gamma>0$.
The asymptotic distribution of the perturbed estimation $\widehat{C}$ admits an explicit expression as detailed below.
Theorem 5.1 (Asymptotic Distribution).
For problem (12), suppose $p\in(1,\infty]$, and $\delta_{n}=\eta n^{-\gamma},\gamma,\eta>0$, then we have that
  • If $0<\gamma<1/2$,
    $n^{\gamma}\left(\widehat{C}-\Sigma^{-1}\right)\Rightarrow-2\eta\Sigma^{-1}\mathbb{E}\left[\frac{\left(\mathrm{sign}(\Sigma^{-1}\bm{x})\odot|\Sigma^{-1}\bm{x}|^{q-1}\right)\bm{x}^{\top}}{\|\Sigma^{-1}\bm{x}\|_{q}^{q-1}}\right]\Sigma^{-1};$
  • If $\gamma=1/2$,
    $\sqrt{n}\left(\widehat{C}-\Sigma^{-1}\right)\Rightarrow\Sigma^{-1}\left(\mathbf{G}-2\eta\mathbb{E}\left[\frac{\left(\mathrm{sign}(\Sigma^{-1}\bm{x})\odot|\Sigma^{-1}\bm{x}|^{q-1}\right)\bm{x}^{\top}}{\|\Sigma^{-1}\bm{x}\|_{q}^{q-1}}\right]\right)\Sigma^{-1};$
  • If $\gamma>1/2$,
    $\sqrt{n}\left(\widehat{C}-\Sigma^{-1}\right)\Rightarrow\Sigma^{-1}\mathbf{G}\Sigma^{-1},$
where $\bm{x}\sim\mathcal{N}(\bm{0},\Sigma)$, $\mathrm{vec}(\mathbf{G})\sim\mathcal{N}(\bm{0},\Lambda)$, and $\Lambda$ is such that
$\mathrm{cov}([\mathbf{G}]_{ij},[\mathbf{G}]_{i^{\prime}j^{\prime}})=\mathrm{cov}([\bm{x}]_{i}[\bm{x}]_{j},[\bm{x}]_{i^{\prime}}[\bm{x}]_{j^{\prime}}).$
The results in Theorem 5.1 have some implications.
Firstly, the proposed perturbed estimation $\widehat{C}$ is a consistent estimator of the precision matrix $\Sigma^{-1}$, i.e., $\widehat{C}\to_{p}\Sigma^{-1}$, when the perturbation magnitude shrinks as the sample size increases. Then, the proposed perturbed estimation has a convergence order of $\sqrt{n}$ when $\gamma\geq 1/2$ and a slower convergence order of $n^{\gamma}$ when $\gamma\in(0,1/2)$.
Secondly, when we focus on cases where $\gamma\geq 1/2$, the perturbed estimation converges in distribution to a normal distribution.
Thirdly, the perturbation magnitude decays fast enough that the bias induced by the adversarial perturbation becomes asymptotically negligible when $\gamma>1/2$. In this regime, the estimation $\widehat{C}$ converges to a zero-mean normal distribution identical to that of the unperturbed inverse sample covariance, i.e., the maximum likelihood estimation from problem (1). This establishes a theoretical threshold: while adversarial training typically introduces a trade-off between robustness and accuracy (Tsipras et al., 2019), accuracy and perturbation are not in conflict when the perturbation order exceeds $1/2$ in our proposed adversarially perturbed precision matrix estimation framework.
Recall that we focus on the surrogate objective function instead when we discuss the $\ell_{\infty}$-perturbed estimation in Section 3, as shown in the following:
$\widetilde{C}\in\arg\inf_{C\succ 0}\left\{-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{x}^{\top}C\bm{x}\right]+2\delta\sum_{j=1}^{d}\sum_{k=1}^{d}\omega_{k}|[C]_{kj}|+\delta^{2}\|C\|_{1,1}\right\}.$ (13)
In contrast to the asymptotic results in Theorem 5.2, the asymptotic distribution of the estimation $\widetilde{C}$ in problem (13) does not admit a simple closed form but can be characterized as the minimizer of a stochastic process, as shown in the following theorem.
Theorem 5.2 (Asymptotic Distribution when $p=\infty$ with Surrogate Objective).
If $\delta_{n}=\eta n^{-\gamma},\eta,\gamma>0$, for the estimation $\widetilde{C}$ obtained from problem (13), we have that
$\sqrt{n}\left(\widetilde{C}-\Sigma^{-1}\right)\Rightarrow\arg\min_{\bm{U}}W(\bm{U}).$
  • If $0<\gamma<1/2$,
    $W(\bm{U})=$ $\frac{1}{2}\mathrm{tr}(\bm{U}\Sigma\bm{U}\Sigma)$
    $+\sqrt{\frac{8}{\pi}}\eta\sum_{j=1}^{d}\sum_{k=1}^{d}[\Sigma]_{kk}\left([\bm{U}]_{jk}\operatorname{sign}([\Sigma^{-1}]_{jk})\mathbb{I}([\Sigma^{-1}]_{jk}\not=0)\right$.
    $\left.+|[\bm{U}]_{jk}|\mathbb{I}([\Sigma^{-1}]_{jk}=0)\right).$
  • If $\gamma=1/2$,
    $W(\bm{U})=$ $-\mathrm{tr}(\bm{U}\mathbf{G})+\frac{1}{2}\mathrm{tr}(\bm{U}\Sigma\bm{U}\Sigma)$
    $+\sqrt{\frac{8}{\pi}}\eta\sum_{j=1}^{d}\sum_{k=1}^{d}[\Sigma]_{kk}\left([\bm{U}]_{jk}\operatorname{sign}([\Sigma^{-1}]_{jk})\mathbb{I}([\Sigma^{-1}]_{jk}\not=0)\right$.
    $\left.+|[\bm{U}]_{jk}|\mathbb{I}([\Sigma^{-1}]_{jk}=0)\right).$
  • If $\gamma>1/2$,
    $W(\bm{U})=-\mathrm{tr}(\bm{U}\mathbf{G})+\frac{1}{2}\mathrm{tr}(\bm{U}\Sigma\bm{U}\Sigma),$
    i.e.,
    $\sqrt{n}\left(\widetilde{C}-\Sigma^{-1}\right)\Rightarrow\Sigma^{-1}\mathbf{G}\Sigma^{-1},$
where $\mathrm{vec}(\mathbf{G})\sim\mathcal{N}(\bm{0},\Lambda)$, and $\Lambda$ is such that
$\mathrm{cov}([\mathbf{G}]_{ij},[\mathbf{G}]_{i^{\prime}j^{\prime}})=\mathrm{cov}([\bm{x}]_{i}[\bm{x}]_{j},[\bm{x}]_{i^{\prime}}[\bm{x}]_{j^{\prime}}).$
Remark 5.3.
Theorem 5.2 provides the theoretical basis for establishing the sparsity recovery property of the estimation $\widetilde{C}$, as formalized in Corollary 3.2
Theorem 5.2 shows that the estimation $\widetilde{C}$ in problem (13) shares the same asymptotic normal distribution of the estimation $\widehat{C}$ in problem (12). However, for $\gamma=1/2$, the asymptotic distribution of $\widetilde{C}$ becomes non-normal. Similar to the standard $\ell_{1}$-regularized precision matrix estimation (Yuan and Lin, 2007), this non-normality helps encourage sparsity in some sense. Note that the asymptotic term $\Sigma]_{kk}$ presented in Theorem 5.2 represents the impact of adversarial perturbation in our surrogate framework (13).
To have a comprehensive understanding and comparison, we conclude some of the key characteristics of the asymptotic distribution of the perturbed estimation in Table 1.
$\in(1,\infty] p=p$$\infty$, surrogate Rate
$\gamma\in(0,1/2)$ constant constant $n^{\gamma}$
$\gamma=1/2$ normality with bias non-normality with bias, sparsity $\sqrt{n}$
$\gamma>1/2$ normality without bias $\sqrt{n}$
Table 1: Asymptotic Distribution of the Perturbed Estimation
6 Numerical Experiments
In this section, we will deliver numerical experiments for the proposed adversarially perturbed precision matrix estimation scheme and compare it with the $\ell_{1}$-regularized precision matrix estimation (Friedman et al., 2008). The experiments are implemented by R.
6.1 Computation
We will focus on the surrogate $\ell_{\infty}$-perturbed matrix estimation framework (9) discussed in Section 3.
Due to the symmetry of the matrix $C$, we can reformulate problem (9) equivalently as follows,
$\inf_{C\succ 0}\left\{-\log\mathrm{det}C+\mathrm{tr}(\bar{A}C)+\|C\|_{1,\Lambda}\right\}$ (14)
where
$\bar{A}=\mathbb{E}_{\mathbb{P}_{n}}[\bm{x}\bm{x}^{\top}],\quad\|C\|_{1,\Lambda}=\sum_{k=1}^{d}\sum_{j=1}^{d}[\lambda]_{kj}\left|[C]_{kj}\right|$
$\lambda]_{kj}=\delta\mathbb{E}_{\mathbb{P}_{n}}\left[\left|[\bm{x}]_{k}\right|\right]+\delta\mathbb{E}_{\mathbb{P}_{n}}\left[\left|[\bm{x}]_{j}\right|\right]+\delta^{2}.$ (15)
The notation $\|C\|_{1,\Lambda}$ denotes the weighted element-wise $\ell_{1}$-norm of the matrix $C$ with the symmetric weight matrix $\Lambda$. The components of the weight matrix $\Lambda$ are dependent on the moment of the observations and the perturbation magnitude as shown in (15). In this way, problem (14) can be considered as the generalized sparse inverse covariance estimation problem discussed in Hsieh et al. (2014). Thus, problem (14) can be solved by the QUIC algorithm proposed by Hsieh et al. (2014). The QUIC algorithm is a second-order algorithm and has been proven to be superlinearly convergent (Hsieh et al., 2014).
In addition, immediately from Theorem 9 in Hsieh et al. (2014), we can conclude that the precision matrix estimation obtained from the $\ell_{\infty}$-perturbed precision matrix estimation proposed in Section 3 is unique.
Corollary 6.1.
There exists a unique minimizer $\widetilde{C}$ for the optimization problem (14) when $\delta>0$.
As the $\ell_{\infty}$-perturbed precision matrix estimation has the weight matrix dependent on the observed samples, the $\ell_{1}$-regularized precision matrix estimation (Friedman et al., 2008) has a constant weight, as shown in the following:
$\inf_{C\succ 0}\left\{-\log\mathrm{det}C+\mathrm{tr}(\bar{A}C)+\lambda\|C\|_{1}\right\}.$ (16)
We also apply the QUIC algorithm (Hsieh et al., 2014) to solve the optimization problem above.
In practice, we do not penalize the diagonals for either method to promote more desirable numerical performance for all experiments in this section.
6.2 Experiments with Synthetic Data
In this subsection, we will carry out the numerical experiments on the following synthetic models, as suggested in Yuan and Lin (2007).
Model 1. An $AR(1)$ model with $c_{i,i}=1$ and $c_{i,i-1}=c_{i-1,i}=0.5$, for i=1,,di=1,...,d.
Model 2. An $AR(2)$ model with $c_{i,i}=1$, $c_{i,i-1}=c_{i-1,i}=0.5$ and $c_{i,i-2}=c_{i-2,i}=0.25$, for i=1,,di=1,...,d.
Model 3. An $AR(3)$ model with $c_{i,i}=1$, $c_{i,i-1}=c_{i-1,i}=0.4$ and $c_{i,i-2}=c_{i-2,i}=c_{i,i-3}=c_{i-3,i}=0.2,$ for i=1,,di=1,...,d.
Model 4. An $AR(4)$ model with $c_{i,i}=1$, $c_{i,i-1}=c_{i-1,i}=0.4$, $c_{i,i-2}=c_{i-2,i}=c_{i,i-3}=c_{i-3,i}=0.2$ and $c_{i,i-4}=c_{i-4,i}=0.1,$ for i=1,,di=1,...,d.
Model 5. Star model with every node connected to the first node, with $c_{i,i}=1$, $c_{1,i}=c_{i,1}=0.2$ and $c_{i,j}=0$ otherwise, for i,j=1,,di,j=1,...,d.
Model 6. Circle model with $c_{ii}=1$, $c_{i,i-1}=c_{i-1,i}=0.5$ and $c_{1,d}=c_{d,1}=0.4$, for i=1,,di=1,...,d.
We may use the following metrics to evaluate the numerical performance in the rest of this section: true negative rate (TNR), true positive rate (TPR), accuracy (ACC), and Matthews Correlation Coefficient (MCC) (Guilford, 1954). The metrics are defined as follows:
$\mathrm{TNR}=\frac{\mathrm{TN}}{\mathrm{TN}+\mathrm{FP}},\quad\mathrm{TPR}=\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FN}},$
$\mathrm{ACC}=\frac{\mathrm{TN}+\mathrm{TP}}{\mathrm{TN}+\mathrm{FP}+\mathrm{TP}+\mathrm{FN}}$
$\mathrm{MCC}=\frac{\mathrm{TP}\times\mathrm{TN}-\mathrm{FP}\times\mathrm{FN}}{\sqrt{(\mathrm{TP}+\mathrm{FP})(\mathrm{TP}+\mathrm{FN})(\mathrm{TN}+\text{FP})(\text{TN}+\text{FN})}},$
where $\mathrm{TP}$, $\mathrm{TN}$, $\mathrm{FP}$, $\mathrm{FN}$ denote the number of true positives, true negatives, false positives, and false negatives, respectively.
For each model, we run simulations by the QUIC algorithm (Hsieh et al., 2014) with dimension $d=20$ and sample sizes $n=\{30,50,70\}$. We consider an equally spaced grid for the perturbation magnitude $\delta$ on $[0.01,0.3]$ in problem (14) and select the regularization parameter $\lambda$ in problem (16) over an equally spaced grid on $[0.01,0.8]$. For each setting, we run two methods 50 times and record the average and standard error.
6.2.1 Sparsity Recovery and Accuracy
We first compare the performance, i.e., sparsity recovery ability and accuracy, based on Kullback–Leibler (KL) loss, TNR, and MCC. The KL loss is defined as follows:
$\mathrm{KL}=-\log|\bar{C}|+\mathrm{tr}(\bar{C}\Sigma)-\left(-\log|\Sigma^{-1}|+d\right),$
where $\Sigma$ is the ground-truth covariance matrix, and $\bar{C}$ is the estimation of the precision matrix.
We visualize the experimental results for the 6 models in Figure 1, Figure 2, and Figure 3. In the figures, “Perturbed” refers to the performance of our proposed method, whereas “Standard” indicates the performance of the standard $\ell_{1}$-regularized method. The x-axis shows the values of the $\delta$ and, $\lambda$ while the y-axis shows the different metrics.
According to the figures, we could conclude that the proposed method (14) has the ability of sparsity recovery since the TNR of the perturbed method can reach or be close to $1$. In addition, the performances of two methods are comparable in terms of accuracy and sparsity recovery ability.
Figure 1: Accuracy and Sparsity Recovery on $AR(1)$ and $AR(2)$
Figure 2: Accuracy and Sparsity Recovery on $AR(3)$ and $AR(4)$
Figure 3: Accuracy and Sparsity Recovery on the Circle model and the Star model
6.2.2 Adversarial Robustness
In addition to the sparsity recovery ability and the accuracy discussed previously, we also evaluate the adversarial robustness of the methods.
In each simulation trial, three independent datasets are generated from the true distribution: a training set $X_{\text{train}}$ of size $n$, a validation set of size $n$, and a test set $X_{\text{test}}$ of size $5n$. For each sample size $n$, we first identify the optimal choice of $\delta$ for the proposed method and $\lambda$ for the standard $\ell_{1}$-regularized method by minimizing the value of $\mathrm{KL}$ on the validation set. We then derive the estimation $\bar{C}_{1}$ from the training data. Next, we generate adversarial examples from the test data based on $\bar{C}_{1}$. The adversarial perturbation is constructed using a gradient-sign method defined as:
$X_{\text{adv}}=X_{\text{test}}+\varepsilon\text{sign}(X_{\text{test}}\bar{C}_{1}),$
where $\varepsilon$ controls the magnitude of the adversarial perturbation. We vary $\varepsilon$ across a grid from 0 to $1$.
Then, we run the two estimation methods based on the perturbed data $X_{\text{adv}}$ and obtain the estimation $\bar{C}_{2}$. The robustness of each method is evaluated by the $\mathrm{KL}$ loss of $\bar{C}_{2}$.
$\text{adversarial KL}=-\log|\bar{C}_{2}|+\mathrm{tr}(\bar{C}_{2}\Sigma)-\left(-\log|\Sigma^{-1}|+d\right)$
A smaller adversarial KL loss indicates that the estimator is more robust to adversarial data perturbation.
The results are visualized in Figure 4, Figure 5, and Figure 6. The x-axis shows the values of $\varepsilon$ while the value of the adversarial KL. The figures demonstrate that the proposed perturbed method is more adversarially robust than the standard $\ell_{1}$-regularized model for most of the scenarios.
Figure 4: Adversarial Robustness on $AR(1)$ and $AR(2)$
Figure 5: Adversarial Robustness on $AR(3)$ and $AR(4)$
Figure 6: Adversarial Robustness on the Circle model and the Star model
6.3 Experiments with Real Data
In this subsection, we run the two estimation methods on the leukemia dataset (Golub et al., 1999) to compare the associated real-world practical performance.
The dataset comprises 7129 gene expression profiles collected from 72 patients, of whom 47 were diagnosed with Acute Lymphoblastic Leukemia (ALL) and 25 with Acute Myeloid Leukemia (AML). We first select the top $d$ genes with the highest marginal variance. The dimension $d$ is varied across the set $\{20,30,40,50,60\}$. The data is centered and scaled to unit variance within the training set and test set, respectively. Following the Linear Discriminant Analysis (LDA) framework, the classification rule relies on the LDA score. For the observation $\bm{x}$, the LDA score is defined by
$\delta_{k}(\bm{x})=\bm{x}^{T}\hat{C}\hat{\bm{\mu}}_{k}-\frac{1}{2}\hat{\bm{\mu}}_{k}^{T}\hat{C}\hat{\bm{\mu}}_{k}+\log\hat{\pi}_{k},\quad k=1,2,$
where $\hat{\pi}_{k}$ represents the proportion of group $k$ in the training set and $\hat{\bm{\mu}}_{k}$ is the empirical mean of the observations in the training set. The classification rule is $\hat{k}(\bm{x})=\arg\max_{k\in\{1,2\}}\delta_{k}(\bm{x})$. The classification performance is dependent on the estimation accuracy of the precision matrix.
The experiment employs a nested cross-validation scheme. In the outer loop, the data is split into 10 stratified folds. For each training run, we additionally carry out an inner stratified 5‑fold cross‑validation to tune the parameters ($\delta$ in the perturbed method (14) and $\lambda$ in the standard $\ell_{1}$-regularized method (16)) through a grid search within $[0.01,1]$, choosing the value that has the best classification accuracy. With the optimal parameter, we train the model on the full training set and then evaluate the model performance on the held‑out test fold. We replicate this whole scheme $50$ times.
To evaluate the performance, we use ACC, MCC, TNR, and TPR. The average value and the standard error of the metrics for 50 replications are recorded in Table 2. Notably, our perturbed method yields consistently higher ACC and MCC, at the cost of a slight decrease in TPR. In conclusion, the proposed method demonstrates superior efficacy in the precision matrix estimation for the leukemia dataset, outperforming the standard $\ell_{1}$-regularized method.
Method ACC MCC TNR TPR
$d=20$ Perturbed 0.868(0.00360) 0.691 (0.01009) 0.627 (0.01014) 0.995 (0.00137)
Standard 0.864 (0.00313) 0.679 (0.00882) 0.611 (0.00951) 0.997 (0.00101)
$d=30$ Perturbed 0.910 (0.00290) 0.792 (0.00806) 0.741(0.00886) 1.000 (0.000500)
Standard 0.902 (0.00301) 0.775 (0.00760) 0.716 (0.00871) 1.000 (0.000000)
$d=40$ Perturbed 0.933 (0.00285) 0.850 (0.00732) 0.809 (0.00840) 1.000 (0.00040)
Standard 0.921 (0.00305) 0.819 (0.00807) 0.771 (0.00928) 1.000 (0.00000)
$d=50$ Perturbed 0.927 (0.00312) 0.833 (0.00841) 0.798 (0.00926) 0.994 (0.00153)
Standard 0.916 (0.00268) 0.808 (0.00722) 0.760 (0.00778) 0.998 (0.000928)
$d=60$ Perturbed 0.929(0.00247) 0.839 (0.00712) 0.806 (0.00727) 0.993 (0.00143)
Standard 0.912 (0.00305) 0.802 (0.00779) 0.757 (0.00800) 0.994 (0.00148)
Table 2: Average Value and Standard Error of the Metrics
7 Discussion
This paper introduces a versatile perturbed estimation framework for the precision matrix and derives the associated regularization effects and asymptotic distributions. Among different perturbation geometries, the $\ell_{2}$- and $\ell_{\infty}$-perturbations are specifically analyzed. Notably, we demonstrate that the proposed framework recovers Wasserstein shrinkage estimation, achieves sparsity recovery, and improves adversarial robustness. In addition to these properties, future directions may include extending our framework to handle complex data structures, including the unknown group structures studied in (Cheng et al., 2025) and the compositional data investigated in (Zhang et al., 2025).
Appendix A Proofs
A.1 Proof of Theorem 2.1
The proof for Theorem 2.1 is based on the following lemma.
Lemma A.1.
The dual of the optimization problem
$\max_{\|\bm{\Delta}\|_{2}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C(\bm{x}+\bm{\Delta})$
is the following optimization problem
$\bm{x}^{\top}C\bm{x}+\inf_{\lambda I-C\succ 0}\left\{\lambda\delta^{2}+\bm{x}^{\top}C(\lambda I-C)^{-1}C\bm{x}\right\}.$
Proof.
The Lagrangian dual of $\max_{\|\bm{\Delta}\|_{2}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C(\bm{x}+\bm{\Delta})$ is as follows:
$\inf_{\lambda\geq 0}\left\{\lambda\delta^{2}+\max_{\bm{\Delta}}\left\{(\bm{x}+\bm{\Delta})^{\top}C(\bm{x}+\bm{\Delta})-\lambda\|\bm{\Delta}\|_{2}^{2}\right\}\right\},$
which is equivalent to
$\bm{x}^{\top}C\bm{x}+\inf_{\lambda\geq 0}\left\{\lambda\delta^{2}+\max_{\bm{\Delta}}\left\{\bm{\Delta}^{\top}(C-\lambda I)\bm{\Delta}+2\bm{x}^{\top}C\bm{\Delta}\right\}\right\}.$
To make the inner optimization problem finite, we should require $\lambda I-C\succ 0$. In this way, the optimal value will be obtained when
$\bm{\Delta}=(\lambda I-C)^{-1}C\bm{x},$
the resulting optimal value is
$\max_{\bm{\Delta}\in\mathbb{R}^{p}}\left\{\bm{\Delta}^{\top}(C-\lambda I)\bm{\Delta}+2\bm{x}^{\top}C\bm{\Delta}\right\}=\bm{x}^{\top}C(\lambda I-C)^{-1}C\bm{x}$
Now we begin to prove Theorem 2.1.
Proof.
It follows from Lemma A.1 that problem (3) is equivalent to the following:
$\inf_{C\succ 0,\lambda I-C\succ 0}\left\{-\log\mathrm{det}C+\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}^{\top}C\bm{x}_{i}+\lambda\delta^{2}+\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}^{\top}C(\lambda I-C)^{-1}C\bm{x}_{i}\right\}.$
By introducing new variables $t_{i}$, we have that
$\inf_{C\succ 0,\lambda I-C\succ 0}\left\{-\log\mathrm{det}C+\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}^{\top}C\bm{x}_{i}+\lambda\delta^{2}+\frac{1}{n}\sum_{i=1}^{n}t_{i}\right\}.$
$\mathrm{s.t.}\quad\bm{x}_{i}^{\top}C(\lambda I-C)^{-1}C\bm{x}_{i}\leq t_{i}\quad\text{for }1\leq i\leq n.$
Due to the Schur complement, we have the following reformulation:
$\inf_{\lambda,\{t_{i}\},C}$ $-\log\mathrm{det}C+\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}^{\top}C\bm{x}_{i}+\lambda\delta^{2}+\frac{1}{n}\sum_{i=1}^{n}t_{i}$
subject to $C\succ 0,\lambda I-C\succ 0$
$\begin{pmatrix}t_{i}&\bm{x}_{i}^{\top}C\\ C\bm{x}_{i}&\lambda I-C\end{pmatrix}\succeq 0,\quad\forall i=1,\dots,n.$
A.2 Proof of Theorem 2.3
Proof.
It remains to show the following identity:
$\mathrm{tr}(C(\lambda I-C)^{-1}C\bar{A})=\lambda^{2}\mathrm{tr}((\lambda I-C)^{-1}\bar{A})-\mathrm{tr}((\lambda I+C)\bar{A}).$
Since a matrix can commute with its own resolvent matrix, we have that
$\mathrm{tr}(C(\lambda I-C)^{-1}C\bar{A})=\mathrm{tr}(C^{2}(\lambda I-C)^{-1}\bar{A}).$
Then, we can conclude that
$\mathrm{tr}(C(\lambda I-C)^{-1}C\bar{A})-\lambda^{2}\mathrm{tr}((\lambda I-C)^{-1}\bar{A})=\mathrm{tr}((C^{2}-\lambda^{2}I)(\lambda I-C)^{-1}\bar{A}).$
Since
$(C^{2}-\lambda^{2}I)=(C-\lambda I)(C+\lambda I),$
the identity has been proven. ∎
A.3 Proof of Proposition 3.1
Proof.
When $p\geq 1$, we have that
$\max_{\|\bm{\Delta}\|_{p}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C(\bm{x}+\bm{\Delta})$ $\leq\max_{\|\bm{\Delta}\|_{\infty}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C(\bm{x}+\bm{\Delta})$
$=\bm{x}^{\top}C\bm{x}+\max_{\|\bm{\Delta}\|_{\infty}\leq\delta}\left\{2\bm{x}^{\top}C\bm{\Delta}+\bm{\Delta}^{\top}C\bm{\Delta}\right\}$
$\leq\bm{x}^{\top}C\bm{x}+\max_{\|\bm{\Delta}\|_{\infty}\leq\delta}2\bm{x}^{\top}C\bm{\Delta}+\max_{\|\bm{\Delta}\|_{\infty}\leq\delta}\bm{\Delta}^{\top}C\bm{\Delta}$
$=\bm{x}^{\top}C\bm{x}+2\delta\|C\bm{x}\|_{1}+\max_{\|\bm{\Delta}\|_{\infty}\leq\delta}\bm{\Delta}^{\top}C\bm{\Delta}$
$\leq\bm{x}^{\top}C\bm{x}+2\delta\|C\bm{x}\|_{1}+\delta^{2}\|C\|_{1,1}$
Notice that
$\mathbb{E}_{\mathbb{P}_{n}}\left[\|C\bm{x}\|_{1}\right]=\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{d}|[C]_{j}\bm{x}_{i}|$ $\leq\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{d}\sum_{k=1}^{d}|[C]_{kj}||[\bm{x}_{i}]_{k}|$
$=\sum_{j=1}^{d}\sum_{k=1}^{d}|[C]_{kj}|\left(\frac{1}{n}\sum_{i=1}^{n}|[\bm{x}_{i}]_{k}|\right).$
We can conclude that the inequality holds. ∎
A.4 Proof of Proposition 4.1
We first state two propositions before we prove Proposition 4.1.
Proposition A.2.
For vector $b\in\mathbb{R}^{d}$, positive definite matrix $Q\in\mathbb{R}^{d\times d}$ and $p\in(1,\infty)$, we have that
$\max_{\|\Delta\|_{p}\leq 1}b^{\top}\Delta+\varepsilon\Delta^{\top}Q\Delta=\|b\|_{q}+\varepsilon\widetilde{\Delta}^{\top}Q\widetilde{\Delta}+o(\varepsilon),$ (17)
as $\varepsilon\to 0$, where $1/p+1/q=1$, and $\widetilde{\Delta}$ is the unique solution to the problem
$\max_{\|\Delta\|_{p}\leq 1}b^{\top}\Delta.$
Proof.
We first prove that the optimization problem (17) has a unique solution when $\varepsilon\to 0$.
The Lagrangian function of problem (17) is
$L_{\varepsilon}(\Delta,\lambda)=b^{\top}\Delta+\varepsilon\Delta^{\top}Q\Delta-\lambda(\|\Delta\|_{p}^{p}-1).$
Since the constraint set $\|\Delta\|_{p}\leq 1$ is convex, it follows from the KKT condition that the solution to the problem, denoted by $\Delta_{\ast}$ and $\lambda_{\ast}$, follows the equations:
$\|\Delta_{\ast}\|_{p}^{p}-1=0,\quad b+\varepsilon 2Q\Delta_{\ast}-\lambda_{\ast}p\|\Delta_{\ast}\|_{p}^{p-1}|\Delta_{\ast}|^{p-1}\odot\operatorname{sign}(\Delta_{\ast})=0.$ (18)
Assume there are two distinct solutions to problem (17) when $\varepsilon\to 0$, denoted by $\Delta_{\ast}^{1}$ and $\Delta_{\ast}^{2}$. Equivalently, both $\Delta_{\ast}^{1}$ and $\Delta_{\ast}^{2}$ satisfy the equations in (18), indicating
$2\varepsilon Q(\Delta_{\ast}^{1}-\Delta_{\ast}^{2})=\lambda_{\ast}p(|\Delta_{\ast}^{2}|^{p-1}\odot\operatorname{sign}(\Delta_{\ast}^{2})-|\Delta_{\ast}^{1}|^{p-1}\odot\operatorname{sign}(\Delta_{\ast}^{1})),$ (19)
$\|\Delta_{\ast}^{1}\|_{p}=\|\Delta_{\ast}^{2}\|_{p}=1,\quad\Delta_{\ast}^{1}\not=\Delta_{\ast}^{2}$
It is easy to see that the left-hand side of (19) goes to $\bm{0}$ when $\varepsilon\to 0$ while the right-hand side of (19) strictly not equal to $\bm{0}$ when $p\in(1,\infty)$ and $\Delta_{\ast}^{1}\not=\Delta_{\ast}^{2}$. This produces a contradiction. Thus, the optimization problem (17) has a unique solution when $\varepsilon\to 0$ and $p\in(1,\infty)$.
Then, since the objective function is convex, continuous, and differentiable w.r.t $\varepsilon$ for fixed $\Delta$, by Danskin’s theorem (Bonnans and Shapiro, 2013), if we denote
$\phi_{b,Q}(\varepsilon)=\max_{\|\Delta\|_{p}\leq 1}b^{\top}\Delta+\varepsilon\Delta^{\top}Q\Delta,\quad\Delta_{b,Q}(\varepsilon)=\arg\max_{\|\Delta\|_{p}\leq 1}b^{\top}\Delta+\varepsilon\Delta^{\top}Q\Delta,$
then we have that
$\frac{d}{d\varepsilon}\phi_{b,Q}(\varepsilon)=\Delta_{b,Q}(\varepsilon)^{\top}Q\Delta_{{b,Q}}(\varepsilon),$
and
$\frac{d}{d\varepsilon}\phi_{b,Q}(\varepsilon)\Big|_{\varepsilon=0}=\widetilde{\Delta}^{\top}Q\widetilde{\Delta}$,
It follows from Taylor’s expansion that
$\phi_{b,Q}(\varepsilon)=\phi_{b,Q}(0)+\varepsilon\widetilde{\Delta}^{\top}Q\widetilde{\Delta}+o(\varepsilon)=\|b\|_{q}+\varepsilon\widetilde{\Delta}^{\top}Q\widetilde{\Delta}+o(\varepsilon),$
as $\varepsilon\to 0$. ∎
Proposition A.3.
For vector $b\in\mathbb{R}^{d}$ with $b]_{i}\not=0$ for i=1,,di=1,...,d, positive definite matrix $Q\in\mathbb{R}^{d\times d}$, we have that
$\max_{\|\Delta\|_{\infty}\leq 1}b^{\top}\Delta+\varepsilon\Delta^{\top}Q\Delta=\|b\|_{1}+\varepsilon\operatorname{sign}(b)^{\top}Q\operatorname{sign}(b),$ (20)
as $\varepsilon\to 0$.
Proof.
When $\varepsilon=0$, the optimization problem $\max_{\|\Delta\|_{\infty}\leq 1}b^{\top}\Delta$ has the unique solution at $\operatorname{sign}(b)$.
When $\varepsilon\not=0$, the objective function is convex and the constraint set is a convex compact set. In this way, the solution to the optimization problem is taken at the boundary of the constraint set, i.e., the solution $\Delta^{\ast}$ satisfies that $\Delta^{\ast}]_{i}=\{-1,1\}$. Since $b]_{i}\not=0$, when $\varepsilon$ is very small, flipping the sign of $\Delta^{\ast}$ will decrease the value of the objective function. Thus, the optimal solution is $\operatorname{sign}(b)$ when $\varepsilon\to 0$. ∎
Now we begin to prove Proposition 4.1.
Proof.
Notice we have that
$\max_{\|\bm{\Delta}\|_{p}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C(\bm{x}+\bm{\Delta})=\bm{x}^{\top}C\bm{x}+\delta\max_{\|\bm{\Delta}\|_{p}\leq 1}\left\{2(C\bm{x})^{\top}\bm{\Delta}+\delta\bm{\Delta}^{\top}C\bm{\Delta}\right\}.$
In the probability space $(\bm{P},C,\mathcal{F})$ for variable $\bm{x}$, for each $\bm{\omega}\in C$, it follows from Proposition A.2 and Proposition A.3 that
$\max_{\|\bm{\Delta}\|_{p}\leq 1}\left\{2(C\bm{x}(\bm{\omega}))^{\top}\bm{\Delta}+\delta\bm{\Delta}^{\top}C\bm{\Delta}\right\}=2\|C\bm{x}(\bm{\omega})\|_{q}+\delta\bm{v}_{\bm{x}(\bm{\omega})}^{\top}C\bm{v}_{\bm{x}(\bm{\omega})}+\mathcal{R}(\bm{\omega}),$
where
$\lim_{\delta\to 0}\frac{\mathcal{R}(\bm{\omega})}{\delta}=0.$
Since we have that
$|\mathcal{R}(\bm{\omega})|$ $=\left|\max_{\|\bm{\Delta}\|_{p}\leq 1}\left\{2(C\bm{x}(\bm{\omega}))^{\top}\bm{\Delta}+\delta\bm{\Delta}^{\top}C\bm{\Delta}\right\}-2\|C\bm{x}(\bm{\omega})\|_{q}+\delta\bm{v}_{\bm{x}(\bm{\omega})}^{\top}C\bm{v}_{\bm{x}(\bm{\omega})}\right|$ (21)
$\leq 4\|C\bm{x}(\bm{\omega})\|_{q}+2\delta\max_{\|y\|_{p}\leq 1,y\in\mathbb{R}^{d}}y^{\top}Cy$
$\leq 4\|C\|_{p\to q}\|\bm{x}(\bm{\omega})\|_{p}+2\delta\|C\|_{p\to q},$
implying that $\mathcal{R}(\bm{\omega})$ is upper bounded by an integral function. Due to the dominated convergence theorem, we have that
$\lim_{\delta\to 0}\frac{\mathbb{E}_{\bm{P}}[\mathcal{R}]}{\delta}=\int_{\bm{\omega}\in C}\lim_{\delta\to 0}\frac{\mathcal{R}(\bm{\omega})}{\delta}d\bm{P}(\bm{\omega})=0.$
A.5 Proof of Theorem 5.1
Proof.
We define
$\Psi_{n}(C)=-\log\det(C)+\mathbb{E}_{\mathbb{P}_{n}}\left[\max_{\|\bm{\Delta}\|_{p}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C(\bm{x}+\bm{\Delta})\right].$
For $C\succ 0$, the function $\Psi_{n}(C)$ is convex since we have the following:
$\Psi_{n}(\alpha C_{1}+(1-\alpha)C_{2})$ (22)
$=$ $-\log\det(\alpha C_{1}+(1-\alpha)C_{2})+\mathbb{E}_{\mathbb{P}_{n}}\left[\max_{\|\bm{\Delta}\|_{p}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}(\alpha C_{1}+(1-\alpha)C_{2})(\bm{x}+\bm{\Delta})\right]$
$\leq$ $-\alpha\log\det(C_{1})-(1-\alpha)\log\det(C_{2})$
$+$ $\alpha\mathbb{E}_{\mathbb{P}_{n}}\left[\max_{\|\bm{\Delta}\|_{p}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C_{1}(\bm{x}+\bm{\Delta})\right]+(1-\alpha)\mathbb{E}_{\mathbb{P}_{n}}\left[\max_{\|\bm{\Delta}\|_{p}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C_{2}(\bm{x}+\bm{\Delta})\right],$
where the second inequality comes from the convexity of the function $\log\det(C)$.
Case 1: $\gamma=1/2$.
It follows from Proposition 4.1 that
$V_{n}(\bm{U})$
$\overset{\vartriangle}{=}$ $n\left(\Psi_{n}(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U})-\Psi(\Sigma^{-1})\right)$
$=$ $-n\log\det(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U})+n\log\det(\Sigma^{-1})+\sqrt{n}\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{x}^{\top}\bm{U}\bm{x}\right]$
$+2\eta\sqrt{n}\left(\mathbb{E}_{\mathbb{P}_{n}}\left[\left\|(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U})\bm{x}\right\|_{q}\right]-\mathbb{E}_{\mathbb{P}_{n}}\left[\left\|\Sigma^{-1}\bm{x}\right\|_{q}\right]\right)$
$+\eta^{2}\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{v}_{\Sigma^{-1}\bm{x}}^{\top}\Sigma^{-1}\bm{v}_{\Sigma^{-1}\bm{x}}\right]-\eta^{2}\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{v}_{(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U})\bm{x}}^{\top}\left(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U}\right)\bm{v}_{(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U})\bm{x}}\right]$
$+o(1).$
It follows from the proof of Theorem 1 in Yuan and Lin (2007) that
$\log\det(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U})+\log\det(\Sigma^{-1})=-\frac{\mathrm{tr}(\bm{U}\Sigma)}{\sqrt{n}}+\frac{1}{2}\frac{\mathrm{tr}(\bm{U}\Sigma\bm{U}\Sigma)}{n}+o\left(\frac{1}{n}\right).$
We can rewrite the term $\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{x}^{\top}\bm{U}\bm{x}\right]$ as $\mathrm{tr}(\bm{U}\bar{A})$, where
$\bar{A}=\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}\bm{x}_{i}^{\top}.$
In this way, we have that
$-n\log\det(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U})+n\log\det(\Sigma^{-1})+\sqrt{n}\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{x}^{\top}\bm{U}\bm{x}\right]$ (23)
$=$ $\mathrm{tr}(\bm{U}\sqrt{n}(\bar{A}-\Sigma))+\frac{1}{2}\mathrm{tr}(\bm{U}\Sigma\bm{U}\Sigma)+o(1)$
$\Rightarrow$ $-\mathrm{tr}(\bm{U}\mathbf{G})+\frac{1}{2}\mathrm{tr}(\bm{U}\Sigma\bm{U}\Sigma),$
where $\mathrm{vec}(\mathbf{G})\sim\mathcal{N}(\bm{0},\Lambda)$, and $\Lambda$ is such that
$\mathrm{cov}([\mathbf{G}]_{ij},[\mathbf{G}]_{i^{\prime}j^{\prime}})=\mathrm{cov}([\bm{x}]_{i}[\bm{x}]_{j},[\bm{x}]_{i^{\prime}}[\bm{x}]_{j^{\prime}}).$
Notice that we also have
$\sqrt{n}\left(\mathbb{E}_{\mathbb{P}_{n}}\left[\left\|(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U})\bm{x}\right\|_{q}\right]-\mathbb{E}_{\mathbb{P}_{n}}\left[\left\|\Sigma^{-1}\bm{x}\right\|_{q}\right]\right)$ (24)
$\to_{p}\mathbb{E}\left[\frac{\left(\mathrm{sign}(\Sigma^{-1}\bm{x})\odot|\Sigma^{-1}\bm{x}|^{q-1}\right)^{\top}\bm{U}\bm{x}}{\|\Sigma^{-1}\bm{x}\|_{q}^{q-1}}\right].$
Thus,
$V_{n}(\bm{U})\Rightarrow V(\bm{U}),$
where
$V(\bm{U})=-\mathrm{tr}(\bm{U}\mathbf{G})+\frac{1}{2}\mathrm{tr}(\bm{U}\Sigma\bm{U}\Sigma)+2\eta\mathbb{E}\left[\frac{\left(\mathrm{sign}(\Sigma^{-1}\bm{x})\odot|\Sigma^{-1}\bm{x}|^{q-1}\right)^{\top}\bm{U}\bm{x}}{\|\Sigma^{-1}\bm{x}\|_{q}^{q-1}}\right].$
We take the gradient of $V(\bm{U})$ w.r.t. $\bm{U}$ and obtain that
$\mathbf{G}+\Sigma\bm{U}^{\ast}\Sigma+2\eta\mathbb{E}\left[\frac{\left(\mathrm{sign}(\Sigma^{-1}\bm{x})\odot|\Sigma^{-1}\bm{x}|^{q-1}\right)\bm{x}^{\top}}{\|\Sigma^{-1}\bm{x}\|_{q}^{q-1}}\right]=0,$
$\bm{U}^{\ast}=\Sigma^{-1}\left(\mathbf{G}-2\eta\mathbb{E}\left[\frac{\left(\mathrm{sign}(\Sigma^{-1}\bm{x})\odot|\Sigma^{-1}\bm{x}|^{q-1}\right)\bm{x}^{\top}}{\|\Sigma^{-1}\bm{x}\|_{q}^{q-1}}\right]\right)\Sigma^{-1}.$
Since $V(\bm{U})$ is strongly convex due to the positive definiteness of $\Sigma$, we have that
$\sqrt{n}\left(\widehat{C}-\Sigma^{-1}\right)\Rightarrow\Sigma^{-1}\left(\mathbf{G}-2\eta\mathbb{E}\left[\frac{\left(\mathrm{sign}(\Sigma^{-1}\bm{x})\odot|\Sigma^{-1}\bm{x}|^{q-1}\right)\bm{x}^{\top}}{\|\Sigma^{-1}\bm{x}\|_{q}^{q-1}}\right]\right)\Sigma^{-1}.$
Case 2: $\gamma>1/2$.
$V_{n}(\bm{U})$
$\overset{\vartriangle}{=}$ $n\left(\Psi_{n}(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U})-\Psi(\Sigma^{-1})\right)$
$=$ $-n\log\det(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U})+n\log\det(\Sigma^{-1})+\sqrt{n}\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{x}^{\top}\bm{U}\bm{x}\right]$
$+2\eta n^{1/2-\gamma}\sqrt{n}\left(\mathbb{E}_{\mathbb{P}_{n}}\left[\left\|(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U})\bm{x}\right\|_{q}\right]-\mathbb{E}_{\mathbb{P}_{n}}[\left\|\Sigma^{-1}\bm{x}\right\|_{q}]\right)$
$+\eta^{2}n^{1-2\gamma}\left(\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{v}_{\Sigma^{-1}\bm{x}}^{\top}\Sigma^{-1}\bm{v}_{\Sigma^{-1}\bm{x}}\right]-\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{v}_{(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U})\bm{x}}^{\top}\left(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U}\right)\bm{v}_{(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U})\bm{x}}\right]\right)$
$+o(1).$
Since $\gamma>1/2$, we have that
$V_{n}(\bm{U})\Rightarrow V(\bm{U}),$
where
$V(\bm{U})=-\mathrm{tr}(\bm{U}\mathbf{G})+\frac{1}{2}\mathrm{tr}(\bm{U}\Sigma\bm{U}\Sigma).$
Then, we have that
$\sqrt{n}\left(\widehat{C}-\Sigma^{-1}\right)\Rightarrow\Sigma^{-1}\mathbf{G}\Sigma^{-1}.$
Case 3: $0<\gamma<1/2$.
$V_{n}(\bm{U})\overset{\vartriangle}{=}$
$n^{2\gamma}\left(\Psi_{n}(\Sigma^{-1}+\frac{1}{n^{\gamma}}\bm{U})-\Psi(\Sigma^{-1})\right)$
$=$ $-n^{2\gamma}\log\det(\Sigma^{-1}+\frac{1}{n^{\gamma}}\bm{U})+n^{2\gamma}\log\det(\Sigma^{-1})+n^{\gamma}\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{x}^{\top}\bm{U}\bm{x}\right]$
$+2\eta n^{\gamma}\left(\mathbb{E}_{\mathbb{P}_{n}}\left[\left\|(\Sigma^{-1}+\frac{1}{n^{\gamma}}\bm{U})\bm{x}\right\|_{q}\right]-\mathbb{E}_{\mathbb{P}_{n}}[\left\|\Sigma^{-1}\bm{x}\right\|_{q}]\right)$
$+\eta^{2}\left(\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{v}_{\Sigma^{-1}\bm{x}}^{\top}\Sigma^{-1}\bm{v}_{\Sigma^{-1}\bm{x}}\right]-\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{v}_{(\Sigma^{-1}+\frac{1}{n^{\gamma}}\bm{U})\bm{x}}^{\top}\left(\Sigma^{-1}+\frac{1}{n^{\gamma}}\bm{U}\right)\bm{v}_{(\Sigma^{-1}+\frac{1}{n^{\gamma}}\bm{U})\bm{x}}\right]\right)$
$+o(1).$
We have that
$n^{2\gamma}\log\det(\Sigma^{-1}+\frac{1}{n^{\gamma}}\bm{U})+n^{2\gamma}\log\det(\Sigma^{-1})+n^{\gamma}\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{x}^{\top}\bm{U}\bm{x}\right]\Rightarrow\frac{1}{2}\mathrm{tr}(\bm{U}\Sigma\bm{U}\Sigma).$
We conclude that
$V_{n}(\bm{U})\Rightarrow V(\bm{U}),$
where
$V(\bm{U})=\frac{1}{2}\mathrm{tr}(\bm{U}\Sigma\bm{U}\Sigma)+2\eta\mathbb{E}\left[\frac{\left(\mathrm{sign}(\Sigma^{-1}\bm{x})\odot|\Sigma^{-1}\bm{x}|^{q-1}\right)^{\top}\bm{U}\bm{x}}{\|\Sigma^{-1}\bm{x}\|_{q}^{q-1}}\right].$
Then, we have that
$\sqrt{n}\left(\widehat{C}-\Sigma^{-1}\right)\Rightarrow-2\eta\Sigma^{-1}\mathbb{E}\left[\frac{\left(\mathrm{sign}(\Sigma^{-1}\bm{x})\odot|\Sigma^{-1}\bm{x}|^{q-1}\right)\bm{x}^{\top}}{\|\Sigma^{-1}\bm{x}\|_{q}^{q-1}}\right]\Sigma^{-1}.$
A.6 Proof of Theorem 5.2
Proof.
We define
$\Phi_{n}(C)=-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{x}^{\top}C\bm{x}\right]+2\delta\sum_{j=1}^{d}\sum_{k=1}^{d}\left(\frac{1}{n}\sum_{i=1}^{n}|[\bm{x}_{i}]_{k}|\right)|[C]_{kj}|+\delta^{2}\|C\|_{1,1}.$
For $C\succ 0$, the function $\Phi_{n}(C)$ is convex.
Case 1: $\gamma=1/2$.
$W_{n}(\bm{U})\overset{\vartriangle}{=}$ $n\left(\Phi_{n}(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U})-\Phi(\Sigma^{-1})\right)$
$=$ $-n\log\det(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U})+n\log\det(\Sigma^{-1})+\sqrt{n}\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{x}^{\top}\bm{U}\bm{x}\right]$
$+2\eta\sqrt{n}\sum_{j=1}^{d}\sum_{k=1}^{d}\left(\frac{1}{n}\sum_{i=1}^{n}|[\bm{x}_{i}]_{k}|\right)\left(|[\Sigma^{-1}]_{kj}+\frac{1}{\sqrt{n}}[\bm{U}]_{kj}|-|[\Sigma^{-1}]_{kj}|\right)$
$+o(1).$
Based on the proof of Theorem 5.1, we can have that
$\sqrt{n}(\widetilde{C}-\Sigma^{-1})\Rightarrow\arg\min_{\bm{U}}W(\bm{U}),$
where
$W(\bm{U})$ $=-\mathrm{tr}(\bm{U}\mathbf{G})+\frac{1}{2}\mathrm{tr}(\bm{U}\Sigma\bm{U}\Sigma)$
$+2\eta\sum_{j=1}^{d}\sum_{k=1}^{d}\mathbb{E}[|[\bm{x}]_{k}|]\left([\bm{U}]_{jk}\operatorname{sign}([\Sigma^{-1}]_{jk})\mathbb{I}([\Sigma^{-1}]_{jk}\not=0)+|[\bm{U}]_{jk}|\mathbb{I}([\Sigma^{-1}]_{jk}=0)\right).$
Case 2: $\gamma>1/2$.
$W(\bm{U})$ $=-\mathrm{tr}(\bm{U}\mathbf{G})+\frac{1}{2}\mathrm{tr}(\bm{U}\Sigma\bm{U}\Sigma).$
Case 3: $0<\gamma<1/2$.
$W(\bm{U})$ $=\frac{1}{2}\mathrm{tr}(\bm{U}\Sigma\bm{U}\Sigma)$
$+2\eta\sum_{j=1}^{d}\sum_{k=1}^{d}\mathbb{E}[|[\bm{x}]_{k}|]\left([\bm{U}]_{jk}\operatorname{sign}([\Sigma^{-1}]_{jk})\mathbb{I}([\Sigma^{-1}]_{jk}\not=0)+|[\bm{U}]_{jk}|\mathbb{I}([\Sigma^{-1}]_{jk}=0)\right).$
Since we have $\bm{x}\sim\mathcal{N}(\bm{0},\Sigma)$, then we have $\bm{x}]_{k}\sim\mathcal{N}(\bm{0},[\Sigma]_{kk})$. In this way,
$\mathbb{E}[|[\bm{x}]_{k}|]=\sqrt{\frac{2[\Sigma]_{kk}}{\pi}}.$

### v2

@@
 This paper proposes an adversarially perturbed precision matrix estimation framework, motivated by recent developments in adversarial training.
-The proposed framework is versatile for the precision matrix problem since, by adapting to different perturbation geometries, the proposed framework can not only recover the existing distributionally robust method but also inspire a novel moment-adaptive approach to precision matrix estimation, proven capable of sparsity recovery and adversarial robustness.
-Notably, the proposed perturbed precision matrix framework is proven to be asymptotically equivalent to regularized precision matrix estimation, and the asymptotic normality can be established accordingly.
-The resulting asymptotic distribution highlights the asymptotic bias introduced by perturbation and identifies conditions under which the perturbed estimation can be unbiased in the asymptotic sense.
-Numerical experiments on both synthetic and real data demonstrate the desirable performance of the proposed adversarially perturbed approach in practice.
+The proposed framework is versatile for the precision matrix problem since, by adapting to different perturbation geometries, the proposed framework can not only recover the existing distributionally robust method but also achieve high-dimensional model selection consistency under the scale-adaptive incoherence condition, which can be viewed as a relaxation of the classic incoherence condition in the heteroscedastic settings.
+Additionally, the proposed perturbed precision matrix estimation framework is asymptotically equivalent to the regularized precision matrix estimation, and the asymptotic normality can be established accordingly, where the asymptotic bias introduced by perturbation is highlighted.
+Numerical experiments demonstrate the desirable practical performance of the proposed adversarially perturbed approach.
 1 Introduction
-The precision matrix, also known as the inverse covariance matrix, can reveal the conditional correlations between pairs of variables given all other variables (Fan et al., 2016).
+The precision matrix, defined as the inverse of the covariance matrix, characterizes the conditional dependence structure of the data, where a zero off-diagonal entry implies that two variables are conditionally independent given all other variables (Fan et al., 2016).
 Estimating the precision matrix is essential in many problems, including neuroimaging (Huang et al., 2010), network analysis (Torri et al., 2018), portfolio optimization (Procacci and Aste, 2022), and genomics (Schäfer and Strimmer, 2005), among others.
-Suppose $\bm{x}_{1},\ldots,\bm{x}_{n}$ are $n$ independent and identically distributed samples drawn from a $d$-dimensional multivariate normal distribution.
+Suppose $\bm{x}^{1},\ldots,\bm{x}^{n}$ are $n$ independent and identically distributed samples drawn from a $d$-dimensional multivariate normal distribution.
 For simplicity, we assume that the population mean is zero and the population covariance matrix $\Sigma$ is positive definite and unknown.
-The standard approach to obtain the precision matrix estimation, i.e., $C=\Sigma^{-1}$, is to identify the minimizer of the negative log-likelihood function:
+The standard approach to obtain the precision matrix estimation, i.e., $\Sigma^{-1}$, is to identify the minimizer of the negative log-likelihood function:
 $\min_{C\succ 0}-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}[\bm{x}^{\top}C\bm{x}],$ (1)
 where $C\succ 0$ means that $C$ is positive definite, $\mathrm{det}C$ denotes the determinant of the matrix $C$, $\mathbb{P}_{n}$ is the empirical distribution.
-When the matrix dimension is small such that $d<n$, the solution to the maximum likelihood estimation problem is the inverse of the sample covariance matrix $\mathbb{E}_{\mathbb{P}_{n}}[\bm{x}\bm{x}^{\top}]$.
-However, if the matrix dimension is large such that $d>n$, the maximum likelihood estimation problem does not admit a solution.
-Numerous studies have been conducted to explore shrinkage and regularization frameworks for precision matrix estimation to achieve estimations which are invertible, well-conditioned, and enjoy desirable statistical properties such as sparsity (Yuan and Lin, 2007; Ledoit and Wolf, 2012; Nguyen et al., 2022).
-We will show later that the proposed adversarially perturbed framework in this paper can recover some of these shrinkage and regularized effects by selecting different types of perturbations.
+When the matrix dimension is small such that $d<n$, the solution to problem (1) is the inverse of the sample covariance matrix $\mathbb{E}_{\mathbb{P}_{n}}[\bm{x}\bm{x}^{\top}]$.
+However, if the matrix dimension is large such that $d>n$, the problem (1) does not admit a solution.
+Numerous studies have been conducted to explore shrinkage and regularization frameworks for precision matrix to achieve estimations which are invertible, well-conditioned, and enjoy desirable statistical properties such as sparsity (Yuan and Lin, 2007; Ledoit and Wolf, 2012; Nguyen et al., 2022).
+We will show later that the proposed adversarially perturbed framework in this paper can recover some of these shrinkage and regularization effects by selecting different types of perturbations.
 Our proposed adversarially perturbed precision matrix framework is inspired by recent developments in trustworthy machine-learning—adversarial training (Goodfellow et al., 2015).
@@
 where $\|\cdot\|$ is the norm of the perturbation $\bm{\Delta}$, and $\delta$ is the magnitude of the perturbation.
-Instead of minimizing the empirical average of the likelihood function, the adversarially perturbed estimation problem (2) seeks for the minimizer of the loss under the most adversarial perturbation $\bm{\Delta}$, which is determined by both $C$ and $\bm{x}$ through the inner maximization problem in (2), subject to the constraint that the norm of the perturbation is bounded.
+Instead of optimizing the empirical average of the likelihood function, the adversarially perturbed estimation problem (2) seeks for the minimizer of the loss under the most adversarial perturbation $\bm{\Delta}$, which is determined by both $C$ and $\bm{x}$ through the inner maximization problem in (2), subject to the constraint that the norm of the perturbation is bounded.
 In this paper, we will focus on the $\ell_{p}$-norm of the perturbation.
@@
 The resulting optimization problem is tractable and yields a nonlinear shrinkage estimator that is well-conditioned, rotation equivalent, and preserves the eigenvalue order (Nguyen et al., 2022).
-When $p=\infty$, by optimizing a convex upper-bound surrogate of the objective function in the outer minimization in problem (2), the resulting estimation can be proven to promote sparsity and adversarial robustness.
-We demonstrate that introducing perturbations is asymptotically equivalent to imposing regularization in precision matrix estimation, provided the perturbation magnitude is small.
-The corresponding first-order and second-order regularization terms are also given.
+When $p=\infty$, we optimize a convex upper-bound surrogate of the objective function in problem (2) to admit tractability.
+The resulting estimation is scale-adaptive and proven to achieve adversarial robustness and high-dimensional model selection consistency.
+The high-dimensional model selection consistency achieved by the $\ell_{1}$-regularized estimation requires the mutual incoherence condition (raskutti2008model; wainwright2009sharp; Ravikumar et al., 2011).
+However, our $\ell_{\infty}$-perturbed precision matrix estimation reduces the upper bound of the mutual incoherence by a factor of the ratio of the maximum to the minimum marginal standard deviation of the variables.
+This property facilitates consistency across a broader range of scenarios, particularly under heteroscedastic settings.
+Then, we demonstrate that introducing perturbations is asymptotically equivalent to imposing regularization in precision matrix estimation, provided the perturbation magnitude is small.
 Based on the regularization effect, the asymptotic distribution of the adversarially perturbed estimation is derived.
 The asymptotic distribution characterizes the bias induced by the inner maximization problem in our proposed framework (2).
-We suppose the perturbation magnitude follows the rule $\delta=\eta/n^{\gamma}$, where $\eta>0$ and $\gamma>0$.
+We suppose the perturbation magnitude follows the rule $\delta=\eta n^{-\gamma}$, where $\eta>0$ and $\gamma>0$.
 Our analysis yields the following two regimes: When $\gamma>1/2$, the asymptotic bias is zero.
@@
 When $\gamma\leq 1/2$, the asymptotic distribution has a non-zero bias.
-Furthermore, specifically for $\ell_{\infty}$-norm perturbations using the upper-bound surrogate of the objective function, the derived asymptotic distribution demonstrates that sparsity recovery is achievable when $\gamma=1/2$.
-The regularization reformulation and asymptotic distribution provide insights into the statistical behavior of both Wasserstein shrinkage estimation and the sparse estimation recovered by our novel framework.
+Furthermore, specifically for $\ell_{\infty}$-perturbed surrogate problem, the derived asymptotic distribution demonstrates that sparsity recovery is achievable when $\gamma=1/2$.
 See a summary table of the associated asymptotic behavior in Table 1, Section 5.
-Numerical experiments are delivered for the case when $p=\infty$ and the surrogate objective is considered.
-We carry out the numerical experiments on both synthetic data and a real-world dataset.
-Firstly, the numerical experiments on the synthetic data validate our theoretical results that the perturbed estimation can recover the sparsity and achieve adversarial robustness.
-Secondly, by adopting linear discriminant analysis based on precision matrix estimation for the real dataset, our method shows more desirable practical performance.
 1.1 Related Work
-1.1.1 Adversarial Training
+Adversarial Training:
 As the modern artificial intelligence and machine learning models grow more powerful and widely used, their reliability becomes increasingly important.
 One major challenge to reliability is that machine learning models can be vulnerable to small changes in input data, which may lead to unreliable outputs (Goodfellow et al., 2015).
-To improve the robustness of the models, adversarial training has been introduced as a principled approach by optimizing against adversarial perturbations posed in the input data (Goodfellow et al., 2015; Madry et al., 2018).
-Different from the traditional empirical risk minimization approach, the resulting adversarial training problem is a min-max robust optimization problem and thus introduces a new statistical estimation paradigm.
-While most of the relevant existing work focus on the vector-valued parameter (Goodfellow et al., 2015; Madry et al., 2018; Javanmard and Soltanolkotabi, 2022; Ribeiro et al., 2023; Xie and Huo, 2024, 2025), our proposed framework provides novel insights on how the adversarial training motivates a new estimation framework for the matrix-valued parameter.
-The proposed framework serves as an alternative for practitioners when the precision matrix estimation with desirable properties such as sparsity and adversarial robustness is needed.
-Adversarial training is closely related to the distributionally robust optimization.
+To improve the model robustness, adversarial training has been introduced as a principled approach by optimizing against perturbations posed in the input data (Goodfellow et al., 2015; Madry et al., 2018).
+Different from the classic empirical risk minimization approach, the resulting adversarial training problem is a min-max robust optimization problem and thus introduces a new statistical estimation paradigm.
+While most of the relevant existing work focus on the vector-valued parameter in Euclidean space (Goodfellow et al., 2015; Madry et al., 2018; Javanmard and Soltanolkotabi, 2022; Ribeiro et al., 2023; Xie and Huo, 2024, 2025), our proposed framework illustrates how adversarial training inspires a novel methodology for estimating matrix-valued parameters within the positive-definite cone.
+To the best of our knowledge, we are the first to formulate the precision matrix estimation problem based on adversarial training.
+Adversarial training is closely related to distributionally robust optimization.
 It is well known that adversarial training can be viewed as a special case of the Wasserstein distributionally robust optimization (Staib and Jegelka, 2017; Gao et al., 2024).
-This paper, however, provides another perspective by demonstrating that, in the context of precision matrix estimation, the Wasserstein distributionally robust optimization is a special case of adversarial training.
-1.1.2 Precision Matrix Estimation
+This paper, however, provides a reverse perspective by demonstrating that, in the context of precision matrix estimation, the Wasserstein distributionally robust optimization is a special case of adversarial training.
+Precision Matrix Estimation:
 Precision matrix estimation has been extensively studied under different settings.
 Classical approaches primarily focus on enforcing sparsity via $\ell_{1}$-regularization or $\ell_{1}$-minimization (Yuan and Lin, 2007; Friedman et al., 2008; Cai et al., 2011).
+The novel framework proposed in this paper is also capable of the sparsity recovery and model selection consistency.
+However, as mentioned in Section 1 and detailed in Section 3.2, the model consistency selection achieved by our framework is based on the scale-adaptive incoherence condition which can be weaker than that for the $\ell_{1}$-regularized estimation, especially under the heteroscedastic settings.
+The reason is that the adaptive incoherence in our framework has a smaller upper bound, potentially relaxing the requirements for model consistency selection when variances differ significantly.
 As applications have broadened, researchers have developed methods robust to data anomalies.
@@
 (2022) addresses distributional shifts.
-Our framework can recover the Wasserstein shrinkage estimation (Nguyen et al., 2022) under $\ell_{2}$-perturbation while being capable of the sparsity recovery achieved by classical $\ell_{1}$-regularized methods (Yuan and Lin, 2007; Friedman et al., 2008).
+From the perspective of robustness, our framework can recover the Wasserstein robust estimation (Nguyen et al., 2022) under $\ell_{2}$-perturbation.
 In addition, by defining the model through an adversarial training lens, we achieve intrinsic adversarial robustness.
-Specifically, as detailed in Section 6.1, under $\ell_{\infty}$-perturbation and a specific surrogate objective, our formulation reduces to a moment-adaptive weighted $\ell_{1}$-regularized precision matrix estimation problem, thereby simultaneously achieving sparse recovery and adversarial robustness.
+Our aforementioned scale adaptivity can also be interpreted as the robustness towards the heteroscedasticity.
 1.2 Notations and Definitions
-$\bm{x}]_{i}$ denotes the $i$th component of vector $\bm{x}$.
-$\operatorname{sign}([\bm{x}]_{i})$ denotes the sign of $\bm{x}]_{i}$ (where we let $\operatorname{sign}(0)=0$), and accordingly we let $\operatorname{sign}(\bm{x})=(\operatorname{sign}([\bm{x}]_{1}),...,\operatorname{sign}([\bm{x}]_{d}))^{\top}$.
-We use $\|\bm{x}\|_{p}$ to denote the $\ell_{p}$-norm of the vector $\bm{z}$, i.e., $\|\bm{x}\|_{p}^{p}=\sum_{j=1}^{d}|x_{j}|^{p},1\leq p<\infty,\|\bm{x}\|_{\infty}=\max_{1\leq j\leq d}|x_{j}|.$ $\odot$ denotes the component-wise product.
+$X_{ij}$ denotes the $(i,j)$-th entry of the matrix $X$.
+$X_{\mathcal{I}\mathcal{J}}$ denotes the submatrix of $X$ defined by the row index set $\mathcal{I}$ and the column index set $\mathcal{J}$.
+$\mathrm{sign}(x)$ denotes the sign of a scalar $x$ with $\mathrm{sign}(0)=0$.
+Accordingly, $\mathrm{sign}(X)$ denotes the element-wise application of the sign function to matrix $X$.
+$\|\bm{x}\|_{p}$ denotes the $\ell_{p}$-norm of the vector $\bm{x}\in\mathbb{R}^{d}$, defined as $\|\bm{x}\|_{p}=(\sum_{j=1}^{d}|x_{j}|^{p})^{1/p}$ for $1\leq p<\infty$, and $\|\bm{x}\|_{\infty}=\max_{1\leq j\leq d}|x_{j}|$.
+$\|X\|_{\infty}$ denotes the induced matrix $\ell_{\infty}$-norm, defined as $\|X\|_{\infty}=\max_{1\leq i\leq d}\sum_{j=1}^{d}|X_{ij}|$.
+$\|X\|_{1,1}$ denotes the element-wise $\ell_{1}$-norm, defined as $\|X\|_{1,1}=\sum_{i=1}^{d}\sum_{j=1}^{d}|[X]_{ij}|$.
+$\|X\|_{\max}$ denotes the maximum element-wise absolute value, defined as $\|X\|_{\max}=\max_{{}_{1\leq i,j\leq d}}|X_{ij}|$.
+$\odot$ denotes the Hadamard product.
+$\otimes$ denotes the Kronecker product.
 $\mathbb{I}(\cdot)$ denotes the indicator function.
@@
 Future work is discussed in Section 7.
-The proofs are relegated to the Supplementary Material whenever possible.
+The proofs are relegated to the appendix whenever possible.
 2 Adversarially $\ell_{2}$-perturbed Precision Matrix Estimation
@@
 $\inf_{C\succ 0}\left\{-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}\left[\max_{\|\bm{\Delta}\|_{2}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C(\bm{x}+\bm{\Delta})\right]\right\},$ (3)
-where the perturbation norm in the constraint set in the inner maximization problem is chosen as $\ell_{2}$ norm.
+where the perturbation norm in the constraint set in the inner maximization problem is chosen as $\ell_{2}$-norm.
 2.1 Tractable Reformulation
@@
 $\inf_{C\succ 0,\lambda I-C\succ 0}\left\{-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}[\bm{x}^{\top}C\bm{x}]+\lambda\delta^{2}+\mathbb{E}_{\mathbb{P}_{n}}[\bm{x}^{\top}C(\lambda I-C)^{-1}C\bm{x}]\right\},$ (4)
-which further admits the following reformulation:
-$\inf_{\lambda,\{t_{i}\},C}$ $-\log\mathrm{det}C+\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}^{\top}C\bm{x}_{i}+\lambda\delta^{2}+\frac{1}{n}\sum_{i=1}^{n}t_{i}$ (5)
+which further admits the following equivalent reformulation:
+$\inf_{\lambda,\{t_{i}\},C}$ $-\log\mathrm{det}C+\frac{1}{n}\sum_{i=1}^{n}{\bm{x}^{i}}^{\top}C\bm{x}^{i}+\lambda\delta^{2}+\frac{1}{n}\sum_{i=1}^{n}t_{i}$ (5)
 $\mathrm{subject\ to}$ $C\succ 0,\lambda I-C\succ 0,$
-$\begin{pmatrix}t_{i}&\bm{x}_{i}^{\top}C\\ C\bm{x}_{i}&\lambda I-C\end{pmatrix}\succeq 0,\quad\forall i=1,\dots,n.$
+$\begin{pmatrix}t_{i}&{\bm{x}^{i}}^{\top}C\\ C\bm{x}_{i}&\lambda I-C\end{pmatrix}\succeq 0,\quad\forall i=1,\dots,n.$
 The reformulation (4) is obtained by deriving the dual of the inner maximization problem in (3), which admits a closed-form solution under $\ell_{2}$-perturbation.
@@
 Problem (5) is a convex semidefinite optimization problem with a log-determinant barrier term in the objective function and linear matrix inequalities in the decision variables $\lambda,\{t_{i}\},C$.
-Such problems fall within the framework of semidefinite (conic) programming and can be solved efficiently by interior-point methods (Boyd and Vandenberghe, 2004; Nemirovski, 2007).
+Such problems fall within the scope of semidefinite (conic) programming and can be solved efficiently by interior-point methods (Boyd and Vandenberghe, 2004; Nemirovski, 2007).
 2.2 Equivalence to the Wasserstein Shrinkage Estimation
@@
 For any fixed $\rho>0$, the Wasserstein distributionally robust estimation problem (6) is equivalent to
-$\inf_{C\succ 0,\lambda I-C\succ 0}\left\{-\log\mathrm{det}C+\lambda\left(\rho^{2}-\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}\bm{x}_{i}^{\top}\right)+\lambda^{2}\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}(\lambda I-C)^{-1}\bm{x}_{i}^{\top}\right\}.$ (7)
+$\inf_{C\succ 0,\lambda I-C\succ 0}\left\{-\log\mathrm{det}C+\lambda\left(\rho^{2}-\frac{1}{n}\sum_{i=1}^{n}\bm{x}^{i}{\bm{x}^{i}}^{\top}\right)+\lambda^{2}\frac{1}{n}\sum_{i=1}^{n}\bm{x}^{i}(\lambda I-C)^{-1}{\bm{x}^{i}}^{\top}\right\}.$ (7)
 The equivalence of our adversarially perturbed estimation problem (3) and distributionally robust trained problem (6) can be built upon problem (4) and (7) based on algebraic calculations.
@@
 (2022), the resulting estimator is invertible, rotation equivalent, preserves the eigenvalue order, and can be obtained by solving a sequential quadratic approximation algorithm.
-Although arriving at the same solution, our proposed perturbed framework (3) adopts a sample-wise perturbation perspective.
-Crucially, this formulation admits a tractable asymptotic expansion, from which the asymptotic distribution is derived, seeing Section 5.
-This framework elucidates the statistical effect of shrinkage and enables statistical inference.
+Although arriving at the same solution, our proposed perturbed framework (3) adopts a sample-wise perturbation perspective and admits the asymptotic expansion, from which the asymptotic distribution is derived, seeing Section 4 and Section 5.
+In this way, our framework also helps facilitate rigorous statistical inference.
 3 Adversarially $\ell_{\infty}$-perturbed Precision Matrix Estimation
@@
 $\inf_{C\succ 0}\left\{-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}\left[\max_{\|\bm{\Delta}\|_{\infty}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C(\bm{x}+\bm{\Delta})\right]\right\}.$ (8)
-3.1 Tractable Surrogate
-Notably, the inner maximization problem in problem (8) is NP-hard (Nemirovski et al., 1999).
+The inner maximization problem in problem (8) is NP-hard (Nemirovski et al., 1999).
 To ensure computational tractability, we replace the NP-hard inner maximization in (8) with a convex upper-bound surrogate.
 Based on the triangle inequality on the perturbation terms, we derive the following formulation:
-$\inf_{C\succ 0}\left\{-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{x}^{\top}C\bm{x}\right]+2\delta\sum_{j=1}^{d}\sum_{k=1}^{d}\omega_{k}|[C]_{kj}|+\delta^{2}\|C\|_{1,1}\right\}$ (9)
+$\inf_{C\succ 0}\left\{-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{x}^{\top}C\bm{x}\right]+2\delta\sum_{j=1}^{d}\sum_{k=1}^{d}\hat{\omega}_{k}|C_{kj}|+\delta^{2}\|C\|_{1,1}\right\},$ (9)
 where
-$\omega_{k}=\mathbb{E}_{\mathbb{P}_{n}}\left[\left|[\bm{x}]_{k}\right|\right].$
-The formulation (9) can be interpreted as a regularized precision matrix estimation with penalties with orders $O(\delta)$ and $O(\delta^{2})$.
-The term $2\delta\sum_{j,k}\omega_{k}|[C]_{kj}|$ scales with $O(\delta)$ and functions as a moment-adaptive weighted $\ell_{1}$-regularizer.
-The weights $\omega_{k}$’s enforce a heterogeneous shrinkage scheme where entries associated with variables of larger average magnitude are penalized more heavily.
+$\hat{\omega}_{k}=\mathbb{E}_{\mathbb{P}_{n}}\left[\left|\bm{x}_{k}\right|\right].$
+The formulation (9) is a regularized precision matrix.
+The term $2\delta\sum_{j,k=1}^{d}\hat{\omega}_{k}|C_{kj}|$ is scale-adaptive weighted $\ell_{1}$-regularizer.
+The weights $\omega_{k}$’s enforce a heterogeneous shrinkage scheme where entries associated with variables of larger scale are penalized more heavily.
 Since high-magnitude features act as high-leverage points for adversarial exploitation, this weighting scheme promotes an estimation that is not only sparse but also intrinsically robust to adversarial manipulation.
@@
 By penalizing the entry-wise magnitude of matrix $C$, this quadratic regularizer serves as a global stability adjustment, ensuring the surrogate remains a rigorous upper bound for the original objective in (8).
-3.1.1 Properties
-We will discuss the properties of the estimation obtained from (9), including the adversarial robustness stated in Proposition 3.1 and the sparsity recovery stated in Corollary 3.2.
+This whole regularization mechanism promotes scale-adaptive model selection and adversarial robustness, which we will be illustrated in Section 3.1 and Section 3.2.
+3.1 Adversarial Robustness
+We demonstrate the adversarial robustness of problem (9) in the following proposition.
 Proposition 3.1 (Adversarial Robustness).
@@
 This property is due to the nature of adversarial training, the motivation behind our proposed problem, and the way we formulate problem (9) as an upper-bound surrogate.
-Corollary 3.2 (Sparsity Recovery).
-Let $\widetilde{C}$ be the minimizer of problem (9).
-Suppose that $\delta=\eta/\sqrt{n}$ with $\eta>0$.
-Then, as $n\to\infty$, the asymptotic distribution of $\widetilde{C}]_{ij}$ has a positive probability mass at 0 whenever $\Sigma^{-1}]_{ij}=0$.
-Remark 3.3.
-Corollary 3.2 follows immediately from Theorem 5.2 in Section 5 and Proposition 3.8 in Xie and Huo (2025).
-Corollary 3.2 implies that for zero entries in the true precision matrix $\Sigma^{-1}$, the asymptotic distribution of the corresponding estimates from (9) has a positive probability at zero.
-We call this property sparsity recovery, which is theoretically guaranteed if the adversarial perturbation order is $1/\sqrt{n}$.
-The sparsity is induced by the first penalty term in (9).
-Notably, while similar to the shrinkage achieved in $\ell_{1}$-regularized precision matrix estimation (Yuan and Lin, 2007; Friedman et al., 2008), this term incorporates the adaptive adjustment for adversarial robustness discussed earlier
+3.2 High-dimensional Consistency
+This subsection will investigate the high-dimensional convergence rate and model selection consistency for the estimation obtained from (9).
+The analysis is based on the primal-dual witness technique innovated by wainwright2009sharp.
+Our analysis extends the primal-dual witness technique to scenarios involving data-dependent and non-uniform penalty parameters.
+We first reformulate problem (9) equivalently as follows
+$\widetilde{C}\in\arg\inf_{C\succ 0}\left\{-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}[\bm{x}^{\top}C\bm{x}]+\sum_{k\neq j}\hat{\lambda}_{kj}\left|C_{kj}\right|\right\},$ (10)
+$\hat{\lambda}_{kj}=\delta\mathbb{E}_{\mathbb{P}_{n}}\left[\left|\bm{x}_{k}\right|\right]+\delta\mathbb{E}_{\mathbb{P}_{n}}\left[\left|\bm{x}_{j}\right|\right]+\delta^{2},$
+where we do not penalize the diagonals to simplify the analysis.
+We then present several notations and necessary assumptions for our subsequent analysis.
+Let the off-diagonal support index set of the ground-truth precision matrix $\Sigma^{-1}$ be $E:=\{(i,j):i\not=j,[\Sigma^{-1}]_{ij}\neq 0,i,j=1,\dots,d\},$ and the augmented support index set with diagonals included is denoted by $S:=E\cup\{(i,i):i=1,\dots,d\}.$ Define the maximum number of nonzeros per row in $\Sigma^{-1}$ with diagonals counted as $s:=\max_{1\leq i\leq d}\#\{j:\ [\Sigma^{-1}]_{ij}\neq 0,j=1,\dots,d\}.$ Let $\omega_{i}^{\ast}:=\mathbb{E}\left[\left|\bm{x}_{i}\right|\right]=\sqrt{2{(\pi)^{-1}}\Sigma_{ii}}$ and define $\lambda^{\ast}_{ij}:=\delta(\omega_{i}^{\ast}+\omega_{j}^{\ast})+\delta^{2},i\neq j,\lambda^{\ast}_{ii}:=0,i,j=1,\dots,d$.
+Let $\Lambda^{\ast}$ be the symmetric matrix with entries $\lambda^{\ast}_{ij}$.
+Let $\sigma_{\max}:=\max_{1\leq i\leq d}\sqrt{\Sigma_{ii}}$, $\sigma_{\min}:=\min_{1\leq i\leq d}\sqrt{\Sigma_{ii}}$.
+Define $\omega_{\max}^{\ast}:=\max_{1\leq i\leq d}\omega_{i}^{\ast}=\sqrt{2{(\pi)^{-1}}}\sigma_{\max},$ $\omega_{\min}^{\ast}:=\min_{1\leq i\leq d}\omega_{i}^{\ast}=\sqrt{2{(\pi)^{-1}}}\sigma_{\min}.$ We also let Hessian take the form $\Gamma=\Sigma\otimes\Sigma$.
+Assumption 3.2 (Covariance Control).
+$\kappa_{\Gamma}:=\|(\Gamma_{SS})^{-1}\|_{\infty}<\infty,\qquad\kappa_{\Sigma}:=\|\Sigma\|_{\infty}<\infty.$
+Assumption 3.3 (Scale-adaptive Incoherence Condition).
+Define $A:=\Gamma_{S^{c}S}(\Gamma_{SS})^{-1},$ and
+$\kappa_{A}:=\|A\|_{\infty}<\infty,\quad\psi^{\ast}:=\max_{e\in S^{c}}\sum_{s\in S}|A_{es}|\frac{\lambda^{\ast}_{s}}{\lambda^{\ast}_{e}}\leq 1-\alpha,\quad\alpha\in(0,1).$ (11)
+Theorem 3.4 (Convergence Rate and Model Selection Consistency).
+Suppose Assumptions 3.23.3 hold. Fix any $\tau>2$ and set
+$\delta=c_{\delta}\sqrt{\frac{\log d}{n}},\qquad c_{\delta}=\frac{\sqrt{\tau}}{\alpha}\max\left\{16C_{2}\sigma_{\max},\ \ \frac{2\sqrt{2\pi}(1+\kappa_{A})C_{1}\sigma_{\max}^{2}}{\sigma_{\min}}\right\},$
+where $C_{1},C_{2}$ are some universal constants.
+Define the constant $B:=C_{1}\sigma_{\max}^{2}\sqrt{\tau}+3\sqrt{2{(\pi)^{-1}}}\sigma_{\max}c_{\delta}+c_{\delta}^{2}.$ Assume the sample size satisfies
+$n\geq\max\left\{2\pi C_{2}^{2}\left(\frac{\sigma_{\max}}{\sigma_{\min}}\right)^{2}\tau,36\kappa_{\Gamma}^{4}\kappa_{\Sigma}^{6}B^{2}s^{2},\left(\frac{12\sqrt{2\pi}(1+\kappa_{A})\kappa_{\Sigma}^{3}\kappa_{\Gamma}^{2}B^{2}}{\alpha\sigma_{\min}c_{\delta}}\right)^{2}s^{2}\right\}\log d.$
+Then, for the solution $\widetilde{C}$ to problem (10), with probability at least $1-4d^{2-\tau}$, the following holds:
+$\|\widetilde{C}-\Sigma^{-1}\|_{\max}\ \leq\ 2\kappa_{\Gamma}B\sqrt{\frac{\log d}{n}},\quad\widetilde{C}_{S^{c}}=0.$
+If the inequality
+$\min_{(i,j)\in E}|[\Sigma^{-1}]_{ij}|\ \geq\ 4\kappa_{\Gamma}B\sqrt{\frac{\log d}{n}}$
+holds, then $\mathrm{sign}(\widetilde{C}_{ij})=\mathrm{sign}([\Sigma^{-1}]_{ij})$ for all $(i,j)\in E$.
+Theorem 3.4 establishes an error rate of order $\sqrt{\log d/n}$, support recovery, and sign consistency.
+While these rates align with classic results for the $\ell_{1}$-regularized estimation (Ravikumar et al., 2011), the conditions required to achieve them are different.
+In the classic $\ell_{1}$-regularized precision matrix estimation, a uniform penalty is applied to all entries.
+The associated model selection consistency is established under the incoherence condition:
+$\mu^{\ast}:=\max_{e\in S^{c}}\sum_{s\in S}|A_{es}|\leq 1-\alpha,\quad\alpha\in(0,1).$
+This condition requires that no null edge is highly correlated with the set of true edges.
+Our perturbed estimation replaces this uniform requirement with the scale-adaptive incoherence condition, as shown in Assumption 3.3.
+The distinction lies in the penalty ratio $\lambda^{*}_{s}/\lambda^{*}_{e}$, which essentially compares the marginal standard deviation of the variables involved in a true edge versus a null edge.
+We claim that the scale-adaptive incoherence condition embedded in our perturbed estimation (10) is superior over the classic one in the $\ell_{1}$-regularized estimation under heteroscedastic settings.
+This can be seen from the following proposition.
+Proposition 3.5.
+Suppose the covariance matrix $\Sigma$ admits the decomposition $\Sigma=DRD$, where $D=\mathrm{diag}(\sqrt{\Sigma_{11}},\dots,\sqrt{\Sigma_{dd}})$. Let $\Gamma^{\dagger}:=R\otimes R$. Define
+$A^{\dagger}:=\Gamma_{S^{c}S}^{\dagger}(\Gamma_{SS}^{\dagger})^{-1}.$
+If $10\delta<\sqrt{2(\pi)^{-1}}\sigma_{\min}$, then
+$\mu^{\ast}=\max_{e\in S^{c}}\sum_{s\in S}|A_{es}|\leq\Big(\frac{\sigma_{\max}}{\sigma_{\min}}\Big)^{2}\|A^{\dagger}\|_{\infty},\quad\psi^{\ast}=\max_{e\in S^{c}}\sum_{s\in S}|A_{es}|\frac{\lambda^{\ast}_{s}}{\lambda^{\ast}_{e}}\leq\frac{21}{20}\Big(\frac{\sigma_{\max}}{\sigma_{\min}}\Big)\|A^{\dagger}\|_{\infty}.$
+Compared to the standard incoherence $\mu^{\ast}$, the scale-adaptive version $\psi^{\ast}$ reduces the incoherence upper bound by a factor of $\sigma_{\max}(\sigma_{\min})^{-1}$, indicating a possible relaxation of the incoherence condition under the heteroscedastic settings.
+The improvement will be more significant as the difference between the maximum marginal standard deviation and the minimum one becomes larger.
 4 Regularization Effect
 We discuss the regularization effect induced by the adversarial perturbation in the proposed framework.
-We characterize this regularization effect mathematically in the following proposition.
+We characterize this regularization effect in the following proposition.
 Proposition 4.1 (Regularization Effect).
-Suppose the random variable $\bm{x}$ follows a distribution $\bm{P}$ with $\mathbb{E}_{\bm{P}}[\|\bm{x}\|_{p}]<\infty$ and $\bm{P}([C\bm{x}]_{i}=0)=0$, $C$ is a positive definite matrix with $\|C\|_{p\to q}<\infty$, and $p\in(1,\infty]$ with $1/p+1/q=1$.
+Suppose the random variable $\bm{x}$ follows a distribution $\bm{P}$ with $\mathbb{E}_{\bm{P}}[\|\bm{x}\|_{p}]<\infty$ and $\bm{P}([C\bm{x}]_{i}=0)=0,i=1,\ldots,d$, $C$ is a positive definite matrix with $\|C\|_{p\to q}<\infty$, and $p\in(1,\infty]$ with $1/p+1/q=1$.
 Then, we have that
@@
 as $\delta\to 0$, where
-$\bm{v}_{C\bm{x}}=\frac{\operatorname{sign}(C\bm{x})\odot|C\bm{x}|^{q-1}}{\|C\bm{x}\|_{q}^{q-1}}.$ (10)
+$\bm{v}_{C\bm{x}}=\frac{\mathrm{sign}(C\bm{x})\odot|C\bm{x}|^{q-1}}{\|C\bm{x}\|_{q}^{q-1}}.$ (12)
 Remark 4.2.
-Since the term $\log\det C$ in the proposed framework (2) is deterministic, Proposition 4.1 focuses on the inner maximization problem in (2) to analyze the induced regularization under $\ell_{p}$-perturbation.
+Since the term $\log\det C$ in the proposed framework (2) is deterministic, Proposition 4.1 focuses on the inner expectation in (2) to analyze the induced regularization under $\ell_{p}$-perturbation.
 Instead of only analyzing on the empirical distribution, we consider a more general distribution $\bm{P}$ satisfying certain regularity conditions.
@@
 $\mathbb{E}_{\bm{P}}\left[\max_{\|\bm{\Delta}\|_{\infty}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C(\bm{x}+\bm{\Delta})\right]$
-$=$ $\mathbb{E}_{\bm{P}}[\bm{x}^{\top}C\bm{x}]+2\delta\mathbb{E}_{\bm{P}}\left[\|C\bm{x}\|_{1}\right]+\delta^{2}\mathbb{E}_{\bm{P}}\left[\operatorname{sign}(C\bm{x})^{\top}C\operatorname{sign}(C\bm{x})\right]+o(\delta^{2}),$
+$=$ $\mathbb{E}_{\bm{P}}[\bm{x}^{\top}C\bm{x}]+2\delta\mathbb{E}_{\bm{P}}\left[\|C\bm{x}\|_{1}\right]+\delta^{2}\mathbb{E}_{\bm{P}}\left[\mathrm{sign}(C\bm{x})^{\top}C\mathrm{sign}(C\bm{x})\right]+o(\delta^{2}),$
 $\mathbb{E}_{\bm{P}}\left[\max_{\|\bm{\Delta}\|_{2}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C(\bm{x}+\bm{\Delta})\right]$
@@
 The second-order regularization term reflects the local curvature of the worst-case loss.
-Specifically, this term corresponds to the mean of the Rayleigh quotient of $C$ evaluated with respect to the vector $C\bm{x}$ (when $p=2$) or $\operatorname{sign}(C\bm{x})$ (when $p=\infty$).
+Specifically, this term corresponds to the mean of the Rayleigh quotient of $C$ evaluated with respect to the vector $C\bm{x}$ (when $p=2$) or $\mathrm{sign}(C\bm{x})$ (when $p=\infty$).
 If $C$ has the eigen-decomposition $C=\sum_{j}\lambda_{j}\bm{u}_{j}\bm{u}_{j}^{\top}$, the associated Rayleigh quotient represents a weighted average of the eigenvalues.
@@
 Similarly, for $p=\infty$:
-$\operatorname{sign}(C\bm{x})^{\top}C\operatorname{sign}(C\bm{x})=\sum_{j}\lambda_{j}(\bm{u}_{j}^{\top}\operatorname{sign}(C\bm{x}))^{2}.$
+$\mathrm{sign}(C\bm{x})^{\top}C\mathrm{sign}(C\bm{x})=\sum_{j}\lambda_{j}(\bm{u}_{j}^{\top}\mathrm{sign}(C\bm{x}))^{2}.$
 Thus, the second-order regularization implies that adversarial perturbation induces a shrinkage effect on the eigenvalues of $C$.
@@
 where $\bm{x}\sim\mathcal{N}(\bm{0},\Sigma)$, $\mathrm{vec}(\mathbf{G})\sim\mathcal{N}(\bm{0},\Lambda)$, and $\Lambda$ is such that
-$\mathrm{cov}([\mathbf{G}]_{ij},[\mathbf{G}]_{i^{\prime}j^{\prime}})=\mathrm{cov}([\bm{x}]_{i}[\bm{x}]_{j},[\bm{x}]_{i^{\prime}}[\bm{x}]_{j^{\prime}}).$
+$\mathrm{cov}([\mathbf{G}]_{ij},[\mathbf{G}]_{i^{\prime}j^{\prime}})=\mathrm{cov}(\bm{x}_{i}\bm{x}_{j},\bm{x}_{i^{\prime}}\bm{x}_{j^{\prime}}).$
 The results in Theorem 5.1 have some implications.
 Firstly, the proposed perturbed estimation $\widehat{C}$ is a consistent estimator of the precision matrix $\Sigma^{-1}$, i.e., $\widehat{C}\to_{p}\Sigma^{-1}$, when the perturbation magnitude shrinks as the sample size increases.
-Then, the proposed perturbed estimation has a convergence order of $\sqrt{n}$ when $\gamma\geq 1/2$ and a slower convergence order of $n^{\gamma}$ when $\gamma\in(0,1/2)$.
+Then, the proposed perturbed estimation has a convergence order of $\sqrt{n}$ when $\gamma\geq 1/2$ and a slower convergence order of $n^{\gamma}$ when $0<\gamma<1/2$.
 Secondly, when we focus on cases where $\gamma\geq 1/2$, the perturbed estimation converges in distribution to a normal distribution.
@@
 Recall that we focus on the surrogate objective function instead when we discuss the $\ell_{\infty}$-perturbed estimation in Section 3, as shown in the following:
-$\widetilde{C}\in\arg\inf_{C\succ 0}\left\{-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{x}^{\top}C\bm{x}\right]+2\delta\sum_{j=1}^{d}\sum_{k=1}^{d}\omega_{k}|[C]_{kj}|+\delta^{2}\|C\|_{1,1}\right\}.$ (13)
+$\widetilde{C}\in\arg\inf_{C\succ 0}\left\{-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{x}^{\top}C\bm{x}\right]+2\delta\sum_{j=1}^{d}\sum_{k=1}^{d}\omega_{k}|C_{kj}|+\delta^{2}\|C\|_{1,1}\right\}.$ (15)
 In contrast to the asymptotic results in Theorem 5.2, the asymptotic distribution of the estimation $\widetilde{C}$ in problem (15) does not admit a simple closed form but can be characterized as the minimizer of a stochastic process, as shown in the following theorem.
@@
     $W(\bm{U})=$ $\frac{1}{2}\mathrm{tr}(\bm{U}\Sigma\bm{U}\Sigma)$
-    $+\sqrt{\frac{8}{\pi}}\eta\sum_{j=1}^{d}\sum_{k=1}^{d}[\Sigma]_{kk}\left([\bm{U}]_{jk}\operatorname{sign}([\Sigma^{-1}]_{jk})\mathbb{I}([\Sigma^{-1}]_{jk}\not=0)\right$.
-    $\left.+|[\bm{U}]_{jk}|\mathbb{I}([\Sigma^{-1}]_{jk}=0)\right).$
+    $+\sqrt{\frac{8}{\pi}}\eta\sum_{j=1}^{d}\sum_{k=1}^{d}\Sigma_{kk}\left(\bm{U}_{jk}\mathrm{sign}([\Sigma^{-1}]_{jk})\mathbb{I}([\Sigma^{-1}]_{jk}\not=0)+|\bm{U}_{jk}|\mathbb{I}([\Sigma^{-1}]_{jk}=0)\right).$
   • If $\gamma=1/2$,
     $W(\bm{U})=$ $-\mathrm{tr}(\bm{U}\mathbf{G})+\frac{1}{2}\mathrm{tr}(\bm{U}\Sigma\bm{U}\Sigma)$
-    $+\sqrt{\frac{8}{\pi}}\eta\sum_{j=1}^{d}\sum_{k=1}^{d}[\Sigma]_{kk}\left([\bm{U}]_{jk}\operatorname{sign}([\Sigma^{-1}]_{jk})\mathbb{I}([\Sigma^{-1}]_{jk}\not=0)\right$.
-    $\left.+|[\bm{U}]_{jk}|\mathbb{I}([\Sigma^{-1}]_{jk}=0)\right).$
+    $+\sqrt{\frac{8}{\pi}}\eta\sum_{j=1}^{d}\sum_{k=1}^{d}\Sigma_{kk}\left(\bm{U}_{jk}\mathrm{sign}([\Sigma^{-1}]_{jk})\mathbb{I}([\Sigma^{-1}]_{jk}\not=0)+|\bm{U}_{jk}|\mathbb{I}([\Sigma^{-1}]_{jk}=0)\right).$
   • If $\gamma>1/2$,
@@
 where $\mathrm{vec}(\mathbf{G})\sim\mathcal{N}(\bm{0},\Lambda)$, and $\Lambda$ is such that
-$\mathrm{cov}([\mathbf{G}]_{ij},[\mathbf{G}]_{i^{\prime}j^{\prime}})=\mathrm{cov}([\bm{x}]_{i}[\bm{x}]_{j},[\bm{x}]_{i^{\prime}}[\bm{x}]_{j^{\prime}}).$
+$\mathrm{cov}([\mathbf{G}]_{ij},[\mathbf{G}]_{i^{\prime}j^{\prime}})=\mathrm{cov}(\bm{x}_{i}\bm{x}_{j},\bm{x}_{i^{\prime}}\bm{x}_{j^{\prime}}).$
+Corollary 5.3 (Sparsity Recovery).
+Let $\widetilde{C}$ be the minimizer of problem (9).
+Suppose that $\delta=\eta n^{-1/2}$ with $\eta>0$.
+Then, as $n\to\infty$, the asymptotic distribution of $\widetilde{C}_{ij}$ has a positive probability mass at 0 whenever $\Sigma^{-1}]_{ij}=0$.
 Remark 5.4.
-Theorem 5.2 provides the theoretical basis for establishing the sparsity recovery property of the estimation $\widetilde{C}$, as formalized in Corollary 3.2
+Corollary 5.3 follows immediately from Theorem 5.2 and Proposition 3.8 in Xie and Huo (2025).
 Theorem 5.2 shows that the estimation $\widetilde{C}$ in problem (15) shares the same asymptotic normal distribution of the estimation $\widehat{C}$ in problem (14).
-However, for $\gamma=1/2$, the asymptotic distribution of $\widetilde{C}$ becomes non-normal.
-Similar to the standard $\ell_{1}$-regularized precision matrix estimation (Yuan and Lin, 2007), this non-normality helps encourage sparsity in some sense.
-Note that the asymptotic term $\Sigma]_{kk}$ presented in Theorem 5.2 represents the impact of adversarial perturbation in our surrogate framework (13).
+However, for $\gamma=1/2$, the asymptotic distribution of $\widetilde{C}$ is non-normal.
+The term $\Sigma_{kk}$ presented in Theorem 5.2 is resulted from the adversarial perturbation in the framework (15), showing the asymptotic effect of the scale-adaptivity and adversarial robustness discussed in Section 3.1 and Section 3.2.
+In addition, our asymptotic analysis demonstrates that the sparsity recovery can be achievable when $\gamma=1/2$, seeing Corollary 5.3.
+Recall the high dimensional finite-sample analysis implies that model selection consistency is obtained when $\delta$ is in the order of $n^{-1/2}$.
+This provides a unification of finite-sample and asymptotic perspectives for $\ell_{\infty}$-perturbed estimation, demonstrating that both regimes demand an equivalent scaling of $\delta\asymp n^{-1/2}$ to guarantee model selection consistency.
 To have a comprehensive understanding and comparison, we conclude some of the key characteristics of the asymptotic distribution of the perturbed estimation in Table 1.
-$\in(1,\infty] p=p$$\infty$, surrogate Rate
+$p\in(1,\infty]$, problem (2) $p=\infty$, surrogate, problem (9) Rate
 $\gamma\in(0,1/2)$ constant constant $n^{\gamma}$
@@
 The experiments are implemented by R.
-6.1 Computation
+The code url:https://github.com/YilingXie27/Perturbed-Precision-Matrix-Estimation.
 We will focus on the surrogate $\ell_{\infty}$-perturbed matrix estimation framework (9) discussed in Section 3.
-Due to the symmetry of the matrix $C$, we can reformulate problem (9) equivalently as follows,
-$\inf_{C\succ 0}\left\{-\log\mathrm{det}C+\mathrm{tr}(\bar{A}C)+\|C\|_{1,\Lambda}\right\}$ (14)
-where
-$\bar{A}=\mathbb{E}_{\mathbb{P}_{n}}[\bm{x}\bm{x}^{\top}],\quad\|C\|_{1,\Lambda}=\sum_{k=1}^{d}\sum_{j=1}^{d}[\lambda]_{kj}\left|[C]_{kj}\right|$
-$\lambda]_{kj}=\delta\mathbb{E}_{\mathbb{P}_{n}}\left[\left|[\bm{x}]_{k}\right|\right]+\delta\mathbb{E}_{\mathbb{P}_{n}}\left[\left|[\bm{x}]_{j}\right|\right]+\delta^{2}.$ (15)
-The notation $\|C\|_{1,\Lambda}$ denotes the weighted element-wise $\ell_{1}$-norm of the matrix $C$ with the symmetric weight matrix $\Lambda$.
-The components of the weight matrix $\Lambda$ are dependent on the moment of the observations and the perturbation magnitude as shown in (15).
-In this way, problem (14) can be considered as the generalized sparse inverse covariance estimation problem discussed in Hsieh et al.
-(2014).
-Thus, problem (14) can be solved by the QUIC algorithm proposed by Hsieh et al.
-(2014).
-The QUIC algorithm is a second-order algorithm and has been proven to be superlinearly convergent (Hsieh et al., 2014).
-In addition, immediately from Theorem 9 in Hsieh et al.
-(2014), we can conclude that the precision matrix estimation obtained from the $\ell_{\infty}$-perturbed precision matrix estimation proposed in Section 3 is unique.
-Corollary 6.1.
-There exists a unique minimizer $\widetilde{C}$ for the optimization problem (14) when $\delta>0$.
-As the $\ell_{\infty}$-perturbed precision matrix estimation has the weight matrix dependent on the observed samples, the $\ell_{1}$-regularized precision matrix estimation (Friedman et al., 2008) has a constant weight, as shown in the following:
+While the $\ell_{\infty}$-perturbed precision matrix estimation has penalties dependent on the observed samples, the $\ell_{1}$-regularized precision matrix estimation (Friedman et al., 2008) has a constant penalty, as shown in the following:
 $\inf_{C\succ 0}\left\{-\log\mathrm{det}C+\mathrm{tr}(\bar{A}C)+\lambda\|C\|_{1}\right\}.$ (16)
-We also apply the QUIC algorithm (Hsieh et al., 2014) to solve the optimization problem above.
 In practice, we do not penalize the diagonals for either method to promote more desirable numerical performance for all experiments in this section.
-6.2 Experiments with Synthetic Data
-In this subsection, we will carry out the numerical experiments on the following synthetic models, as suggested in Yuan and Lin (2007).
-Model 1. An $AR(1)$ model with $c_{i,i}=1$ and $c_{i,i-1}=c_{i-1,i}=0.5$, for i=1,,di=1,...,d.
-Model 2. An $AR(2)$ model with $c_{i,i}=1$, $c_{i,i-1}=c_{i-1,i}=0.5$ and $c_{i,i-2}=c_{i-2,i}=0.25$, for i=1,,di=1,...,d.
-Model 3. An $AR(3)$ model with $c_{i,i}=1$, $c_{i,i-1}=c_{i-1,i}=0.4$ and $c_{i,i-2}=c_{i-2,i}=c_{i,i-3}=c_{i-3,i}=0.2,$ for i=1,,di=1,...,d.
-Model 4. An $AR(4)$ model with $c_{i,i}=1$, $c_{i,i-1}=c_{i-1,i}=0.4$, $c_{i,i-2}=c_{i-2,i}=c_{i,i-3}=c_{i-3,i}=0.2$ and $c_{i,i-4}=c_{i-4,i}=0.1,$ for i=1,,di=1,...,d.
-Model 5. Star model with every node connected to the first node, with $c_{i,i}=1$, $c_{1,i}=c_{i,1}=0.2$ and $c_{i,j}=0$ otherwise, for i,j=1,,di,j=1,...,d.
-Model 6. Circle model with $c_{ii}=1$, $c_{i,i-1}=c_{i-1,i}=0.5$ and $c_{1,d}=c_{d,1}=0.4$, for i=1,,di=1,...,d.
-We may use the following metrics to evaluate the numerical performance in the rest of this section: true negative rate (TNR), true positive rate (TPR), accuracy (ACC), and Matthews Correlation Coefficient (MCC) (Guilford, 1954).
+We may use the following metrics to evaluate the numerical performance : true negative rate (TNR), true positive rate (TPR), accuracy (ACC), and Matthews Correlation Coefficient (MCC) (Guilford, 1954).
 The metrics are defined as follows:
 $\mathrm{TNR}=\frac{\mathrm{TN}}{\mathrm{TN}+\mathrm{FP}},\quad\mathrm{TPR}=\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FN}},$
-$\mathrm{ACC}=\frac{\mathrm{TN}+\mathrm{TP}}{\mathrm{TN}+\mathrm{FP}+\mathrm{TP}+\mathrm{FN}}$
+$\mathrm{ACC}=\frac{\mathrm{TN}+\mathrm{TP}}{\mathrm{TN}+\mathrm{FP}+\mathrm{TP}+\mathrm{FN}},$
 $\mathrm{MCC}=\frac{\mathrm{TP}\times\mathrm{TN}-\mathrm{FP}\times\mathrm{FN}}{\sqrt{(\mathrm{TP}+\mathrm{FP})(\mathrm{TP}+\mathrm{FN})(\mathrm{TN}+\text{FP})(\text{TN}+\text{FN})}},$
 where $\mathrm{TP}$, $\mathrm{TN}$, $\mathrm{FP}$, $\mathrm{FN}$ denote the number of true positives, true negatives, false positives, and false negatives, respectively.
-For each model, we run simulations by the QUIC algorithm (Hsieh et al., 2014) with dimension $d=20$ and sample sizes $n=\{30,50,70\}$.
-We consider an equally spaced grid for the perturbation magnitude $\delta$ on $[0.01,0.3]$ in problem (14) and select the regularization parameter $\lambda$ in problem (16) over an equally spaced grid on $[0.01,0.8]$.
-For each setting, we run two methods 50 times and record the average and standard error.
-6.2.1 Sparsity Recovery and Accuracy
-We first compare the performance, i.e., sparsity recovery ability and accuracy, based on Kullback–Leibler (KL) loss, TNR, and MCC. The KL loss is defined as follows:
-$\mathrm{KL}=-\log|\bar{C}|+\mathrm{tr}(\bar{C}\Sigma)-\left(-\log|\Sigma^{-1}|+d\right),$
-where $\Sigma$ is the ground-truth covariance matrix, and $\bar{C}$ is the estimation of the precision matrix.
-We visualize the experimental results for the 6 models in Figure 1, Figure 2, and Figure 3.
-In the figures, “Perturbed” refers to the performance of our proposed method, whereas “Standard” indicates the performance of the standard $\ell_{1}$-regularized method.
-The x-axis shows the values of the $\delta$ and, $\lambda$ while the y-axis shows the different metrics.
-According to the figures, we could conclude that the proposed method (14) has the ability of sparsity recovery since the TNR of the perturbed method can reach or be close to $1$.
-In addition, the performances of two methods are comparable in terms of accuracy and sparsity recovery ability.
-Figure 1: Accuracy and Sparsity Recovery on $AR(1)$ and $AR(2)$
-Figure 2: Accuracy and Sparsity Recovery on $AR(3)$ and $AR(4)$
-Figure 3: Accuracy and Sparsity Recovery on the Circle model and the Star model
-6.2.2 Adversarial Robustness
-In addition to the sparsity recovery ability and the accuracy discussed previously, we also evaluate the adversarial robustness of the methods.
-In each simulation trial, three independent datasets are generated from the true distribution: a training set $X_{\text{train}}$ of size $n$, a validation set of size $n$, and a test set $X_{\text{test}}$ of size $5n$.
-For each sample size $n$, we first identify the optimal choice of $\delta$ for the proposed method and $\lambda$ for the standard $\ell_{1}$-regularized method by minimizing the value of $\mathrm{KL}$ on the validation set.
-We then derive the estimation $\bar{C}_{1}$ from the training data.
-Next, we generate adversarial examples from the test data based on $\bar{C}_{1}$.
-The adversarial perturbation is constructed using a gradient-sign method defined as:
-$X_{\text{adv}}=X_{\text{test}}+\varepsilon\text{sign}(X_{\text{test}}\bar{C}_{1}),$
-where $\varepsilon$ controls the magnitude of the adversarial perturbation. We vary $\varepsilon$ across a grid from 0 to $1$.
-Then, we run the two estimation methods based on the perturbed data $X_{\text{adv}}$ and obtain the estimation $\bar{C}_{2}$.
-The robustness of each method is evaluated by the $\mathrm{KL}$ loss of $\bar{C}_{2}$.
-$\text{adversarial KL}=-\log|\bar{C}_{2}|+\mathrm{tr}(\bar{C}_{2}\Sigma)-\left(-\log|\Sigma^{-1}|+d\right)$
-A smaller adversarial KL loss indicates that the estimator is more robust to adversarial data perturbation.
-The results are visualized in Figure 4, Figure 5, and Figure 6.
-The x-axis shows the values of $\varepsilon$ while the value of the adversarial KL.
-The figures demonstrate that the proposed perturbed method is more adversarially robust than the standard $\ell_{1}$-regularized model for most of the scenarios.
-Figure 4: Adversarial Robustness on $AR(1)$ and $AR(2)$
-Figure 5: Adversarial Robustness on $AR(3)$ and $AR(4)$
-Figure 6: Adversarial Robustness on the Circle model and the Star model
+6.1 Experiments with Synthetic Data
+We will carry out the numerical experiments on the following synthetic models.
+Model 1. An $AR(2)$ model with $c_{i,i}=1$, $c_{i,i-1}=c_{i-1,i}=0.5$ and $c_{i,i-2}=c_{i-2,i}=0.25$, for $i=1,\ldots,d$.
+Model 2. An $AR(3)$ model with $c_{i,i}=1$, $c_{i,i-1}=c_{i-1,i}=0.4$ and $c_{i,i-2}=c_{i-2,i}=c_{i,i-3}=c_{i-3,i}=0.2,$ for $i=1,\ldots,d$.
+Model 3. An $AR(4)$ model with $c_{i,i}=1$, $c_{i,i-1}=c_{i-1,i}=0.4$, $c_{i,i-2}=c_{i-2,i}=c_{i,i-3}=c_{i-3,i}=0.2$ and $c_{i,i-4}=c_{i-4,i}=0.1,$ for $i=1,\ldots,d$.
+Model 4. Star model with every node connected to the first node, with $c_{i,i}=1$, $c_{1,i}=c_{i,1}=0.2$ and $c_{i,j}=0$ otherwise, for $i,j=1,\ldots,d$.
+Model 5. Circle model with $c_{i,i}=1$, $c_{i,i-1}=c_{i-1,i}=0.5$ and $c_{1,d}=c_{d,1}=0.4$, for $i=1,\ldots,d$.
+We inject heteroskedastic marginal variances by applying a diagonal rescaling into the models above to illustrate the superiority of the perturbed method under heterogeneous variance settings, as discussed in Section 3.2.
+Define a diagonal scaling matrix $S=\mathrm{diag}(s_{1},\ldots,s_{d})$ with $s_{i}=10$ for $i=1,\ldots,5$ and $s_{i}=1$ for $i=6,\ldots,d$.
+We then set the data-generating covariance to $\Sigma_{S}=S\Sigma S,$.
+Because $D$ is diagonal, the sparsity pattern of the corresponding precision matrix is unchanged, while the first five variables have substantially larger scales.
+For each model, we run simulations by the QUIC algorithm (Hsieh et al., 2014) with dimension $d=30$ and sample sizes $n=\{20,30,40\}$.
+We consider equally spaced grids for the perturbation magnitude $\delta$ in problem (9) and the regularization parameter $\lambda$ in problem (16) over an equally spaced grid on $[0.01,1]$.
+The value of $\delta$ and $\lambda$ is chosen based on the Bayesian Information Criterion.
+For each setting, we run two methods 100 times and record the average and standard error.
+We report the results of the perturbed method and the $\ell_{1}$-regularized method in Tables 26.
+Since our primary interest lies in heteroscedastic settings, we additionally report the performance of the $\ell_{1}$-regularized method applied to standardized data.
+Across the majority of scenarios, the perturbed method attains higher ACC and MCC.
+This observation is consistent with our theoretical results on improved model selection consistency under heteroscedastic models.
+$n$ Method ACC MCC TNR TPR
+2020 Perturbed 0.876 (0.000) 0.232 (0.004) 0.994 (0.000) 0.095 (0.002)
+$\ell_{1}$ 0.233 (0.001) 0.036 (0.004) 0.132 (0.001) 0.903 (0.004)
+$\ell_{1}$ std 0.869 (0.000) 0.016 (0.005) 0.999 (0.000) 0.006 (0.002)
+$30$ Perturbed 0.877 (0.000) 0.243 (0.003) 0.994 (0.000) 0.102 (0.002)
+$\ell_{1}$ 0.698 (0.002) -0.002 (0.003) 0.769 (0.002) 0.229 (0.005)
+$\ell_{1}$ std 0.871 (0.001) 0.051 (0.011) 0.998 (0.001) 0.026 (0.007)
+$40$ Perturbed 0.879 (0.000) 0.263 (0.004) 0.994 (0.001) 0.115 (0.003)
+$\ell_{1}$ 0.709 (0.001) 0.010 (0.004) 0.781 (0.002) 0.232 (0.005)
+$\ell_{1}$ std 0.875 (0.001) 0.148 (0.018) 0.993 (0.002) 0.096 (0.014)
+Table 2: Average Value and Standard Error of the Metrics for $AR(2)$
+$n$ Method ACC MCC TNR TPR
+2020 Perturbed 0.818 (0.000) 0.218 (0.003) 0.998 (0.000) 0.069 (0.002)
+$\ell_{1}$ 0.318 (0.012) 0.000 (0.005) 0.203 (0.020) 0.796 (0.022)
+$\ell_{1}$ std 0.807 (0.000) 0.005 (0.002) 1.000 (0.000) 0.001 (0.001)
+$30$ Perturbed 0.820 (0.000) 0.228 (0.003) 0.998 (0.000) 0.074 (0.002)
+$\ell_{1}$ 0.662 (0.001) -0.047 (0.002) 0.780 (0.002) 0.171 (0.002)
+$\ell_{1}$ std 0.808 (0.000) 0.027 (0.007) 0.999 (0.000) 0.009 (0.003)
+$40$ Perturbed 0.819 (0.000) 0.223 (0.003) 0.998 (0.000) 0.072 (0.002)
+$\ell_{1}$ 0.672 (0.001) -0.035 (0.003) 0.791 (0.002) 0.173 (0.003)
+$\ell_{1}$ std 0.813 (0.001) 0.085 (0.014) 0.996 (0.001) 0.045 (0.009)
+Table 3: Average Value and Standard Error of the Metrics for $AR(3)$
+$n$ Method ACC MCC TNR TPR
+2020 Perturbed 0.761 (0.000) 0.203 (0.003) 0.999 (0.000) 0.058 (0.001)
+$\ell_{1}$ 0.407 (0.014) -0.029 (0.005) 0.324 (0.029) 0.650 (0.032)
+$\ell_{1}$ std 0.747 (0.000) 0.005 (0.002) 1.000 (0.000) 0.001 (0.000)
+$30$ Perturbed 0.761 (0.000) 0.199 (0.002) 1.000 (0.000) 0.054 (0.001)
+$\ell_{1}$ 0.618 (0.001) -0.078 (0.002) 0.776 (0.002) 0.151 (0.001)
+$\ell_{1}$ std 0.748 (0.000) 0.012 (0.004) 1.000 (0.000) 0.002 (0.001)
+$40$ Perturbed 0.760 (0.000) 0.195 (0.002) 1.000 (0.000) 0.053 (0.001)
+$\ell_{1}$ 0.631 (0.001) -0.059 (0.002) 0.793 (0.002) 0.154 (0.001)
+$\ell_{1}$ std 0.749 (0.001) 0.031 (0.007) 0.999 (0.001) 0.011 (0.004)
+Table 4: Average Value and Standard Error of the Metrics for $AR(4)$
+$n$ Method ACC MCC TNR TPR
+2020 Perturbed 0.866 (0.008) 0.534 (0.016) 0.861 (0.008) 0.929 (0.009)
+$\ell_{1}$ 0.837 (0.009) 0.496 (0.006) 0.826 (0.010) 0.989 (0.002)
+$\ell_{1}$ std 0.641 (0.009) 0.157 (0.006) 0.639 (0.010) 0.666 (0.012)
+$30$ Perturbed 0.886 (0.006) 0.581 (0.011) 0.879 (0.006) 0.981 (0.005)
+$\ell_{1}$ 0.861 (0.001) 0.524 (0.002) 0.851 (0.001) 0.996 (0.001)
+$\ell_{1}$ std 0.644 (0.003) 0.208 (0.004) 0.635 (0.003) 0.772 (0.006)
+$40$ Perturbed 0.902 (0.004) 0.611 (0.007) 0.896 (0.004) 0.997 (0.001)
+$\ell_{1}$ 0.873 (0.001) 0.546 (0.002) 0.864 (0.001) 0.999 (0.000)
+$\ell_{1}$ std 0.628 (0.002) 0.236 (0.003) 0.612 (0.002) 0.854 (0.006)
+Table 5: Average Value and Standard Error of the Metrics for the Star model
+$n$ Method ACC MCC TNR TPR
+2020 Perturbed 0.622 (0.014) 0.310 (0.008) 0.594 (0.015) 0.993 (0.002)
+$\ell_{1}$ 0.398 (0.015) 0.181 (0.007) 0.355 (0.016) 0.980 (0.002)
+$\ell_{1}$ std 0.633 (0.012) 0.320 (0.009) 0.606 (0.013) 0.995 (0.001)
+$30$ Perturbed 0.706 (0.004) 0.363 (0.004) 0.684 (0.005) 1.000 (0.000)
+$\ell_{1}$ 0.587 (0.005) 0.283 (0.003) 0.557 (0.005) 0.998 (0.001)
+$\ell_{1}$ std 0.713 (0.011) 0.381 (0.008) 0.692 (0.012) 0.999 (0.000)
+$40$ Perturbed 0.707 (0.003) 0.363 (0.003) 0.685 (0.004) 1.000 (0.000)
+$\ell_{1}$ 0.606 (0.004) 0.294 (0.002) 0.577 (0.004) 1.000 (0.000)
+$\ell_{1}$ std 0.666 (0.011) 0.346 (0.009) 0.642 (0.012) 1.000 (0.000)
+Table 6: Average Value and Standard Error of the Metrics for the Circle model
 6.2 Experiments with Real Data
@@
 The dimension $d$ is varied across the set $\{20,30,40,50,60\}$.
-The data is centered and scaled to unit variance within the training set and test set, respectively.
+The data is then centered and standardized.
 Following the Linear Discriminant Analysis (LDA) framework, the classification rule relies on the LDA score.
@@
 In the outer loop, the data is split into 10 stratified folds.
-For each training run, we additionally carry out an inner stratified 5‑fold cross‑validation to tune the parameters ($\delta$ in the perturbed method (14) and $\lambda$ in the standard $\ell_{1}$-regularized method (16)) through a grid search within $[0.01,1]$, choosing the value that has the best classification accuracy.
+For each training run, we additionally carry out an inner stratified 5‑fold cross‑validation to tune the parameters, $\delta$ in the perturbed method (9) and $\lambda$ in the standard $\ell_{1}$-regularized method (16), through a grid search within $[0.01,1]$, choosing the value that has the best MCC.
 With the optimal parameter, we train the model on the full training set and then evaluate the model performance on the held‑out test fold.
-We replicate this whole scheme $50$ times.
-To evaluate the performance, we use ACC, MCC, TNR, and TPR.
-The average value and the standard error of the metrics for 50 replications are recorded in Table 2.
-Notably, our perturbed method yields consistently higher ACC and MCC, at the cost of a slight decrease in TPR.
-In conclusion, the proposed method demonstrates superior efficacy in the precision matrix estimation for the leukemia dataset, outperforming the standard $\ell_{1}$-regularized method.
-Method ACC MCC TNR TPR
-$d=20$ Perturbed 0.868(0.00360) 0.691 (0.01009) 0.627 (0.01014) 0.995 (0.00137)
-Standard 0.864 (0.00313) 0.679 (0.00882) 0.611 (0.00951) 0.997 (0.00101)
-$d=30$ Perturbed 0.910 (0.00290) 0.792 (0.00806) 0.741(0.00886) 1.000 (0.000500)
-Standard 0.902 (0.00301) 0.775 (0.00760) 0.716 (0.00871) 1.000 (0.000000)
-$d=40$ Perturbed 0.933 (0.00285) 0.850 (0.00732) 0.809 (0.00840) 1.000 (0.00040)
-Standard 0.921 (0.00305) 0.819 (0.00807) 0.771 (0.00928) 1.000 (0.00000)
-$d=50$ Perturbed 0.927 (0.00312) 0.833 (0.00841) 0.798 (0.00926) 0.994 (0.00153)
-Standard 0.916 (0.00268) 0.808 (0.00722) 0.760 (0.00778) 0.998 (0.000928)
-$d=60$ Perturbed 0.929(0.00247) 0.839 (0.00712) 0.806 (0.00727) 0.993 (0.00143)
-Standard 0.912 (0.00305) 0.802 (0.00779) 0.757 (0.00800) 0.994 (0.00148)
-Table 2: Average Value and Standard Error of the Metrics
+We replicate this whole scheme $100$ times.
+To evaluate performance, we report ACC, MCC, TNR, and TPR.
+We treat AML as the positive class and ALL as the negative class, thus TPR corresponds to sensitivity for AML and TNR corresponds to specificity for ALL.
+The average metrics and their standard errors are computed across the 100 replications, where the standard error is calculated from the variability of the replicate-level means.
+Notably, across the considered values of $d$, the perturbed method generally achieves higher ACC, MCC, and TPR.
+Overall, these results suggest that the proposed perturbed estimator can provide improved precision matrix estimates for this dataset, translating into better LDA classification performance compared with the $\ell_{1}$-regularized approach.
+$d$ Method ACC MCC TNR TPR
+2020 Perturbed 0.866 (0.002) 0.682 (0.005) 0.997 (0.001) 0.618 (0.005)
+$\ell_{1}$ 0.863 (0.002) 0.670 (0.006) 0.997 (0.001) 0.607 (0.006)
+$30$ Perturbed 0.921 (0.002) 0.822 (0.005) 1.000 (0.000) 0.771 (0.006)
+$\ell_{1}$ 0.918 (0.002) 0.817 (0.004) 1.000 (0.000) 0.764 (0.005)
+$40$ Perturbed 0.938 (0.002) 0.863 (0.004) 0.998 (0.001) 0.823 (0.004)
+$\ell_{1}$ 0.937 (0.002) 0.860 (0.004) 0.998 (0.001) 0.820 (0.005)
+$50$ Perturbed 0.944 (0.001) 0.878 (0.003) 0.997 (0.001) 0.844 (0.004)
+$\ell_{1}$ 0.939 (0.001) 0.866 (0.004) 0.998 (0.001) 0.826 (0.004)
+$60$ Perturbed 0.938 (0.001) 0.864 (0.004) 0.990 (0.001) 0.838 (0.004)
+$\ell_{1}$ 0.935 (0.001) 0.859 (0.003) 0.993 (0.001) 0.825 (0.004)
+Table 7: Average Value and Standard Error of the Metrics for leukemia dataset
 7 Discussion
-This paper introduces a versatile perturbed estimation framework for the precision matrix and derives the associated regularization effects and asymptotic distributions.
-Among different perturbation geometries, the $\ell_{2}$- and $\ell_{\infty}$-perturbations are specifically analyzed.
-Notably, we demonstrate that the proposed framework recovers Wasserstein shrinkage estimation, achieves sparsity recovery, and improves adversarial robustness.
-In addition to these properties, future directions may include extending our framework to handle complex data structures, including the unknown group structures studied in (Cheng et al., 2025) and the compositional data investigated in (Zhang et al., 2025).
-Appendix A Proofs
-A.1 Proof of Theorem 2.1
-The proof for Theorem 2.1 is based on the following lemma.
+In addition to robustness and model selection consistency, future directions may include extending our framework to handle more complex data structures, including the unknown group structures studied in (Cheng et al., 2025) and the compositional data investigated in (Zhang et al., 2025).
+While the adversarial perturbations investigated in our framework are sample-wise, as defined in (2), an intriguing direction for future research would be to extend this to feature-wise perturbations.
+In the linear regression, xu2008robust has shown that robust linear regression framework under the feature-wise perturbation is equivalent to LASSO.
+Inspired by this fact, the computational and statistical explorations of feature-wise perturbed precision matrix estimation is of much interest.
+Appendix A Proof of Theorem 2.1
+The proof for Theorem 2.1 is based on the following lemmas.
 Lemma A.1.
@@
 $\begin{pmatrix}t_{i}&\bm{x}_{i}^{\top}C\\ C\bm{x}_{i}&\lambda I-C\end{pmatrix}\succeq 0,\quad\forall i=1,\dots,n.$
-A.2 Proof of Theorem 2.3
+Appendix B Proof of Theorem 2.3
 Proof.
@@
 the identity has been proven. ∎
-A.3 Proof of Proposition 3.1
+Appendix C Proof of Proposition 3.1
 Proof.
@@
 Notice that
-$\mathbb{E}_{\mathbb{P}_{n}}\left[\|C\bm{x}\|_{1}\right]=\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{d}|[C]_{j}\bm{x}_{i}|$ $\leq\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{d}\sum_{k=1}^{d}|[C]_{kj}||[\bm{x}_{i}]_{k}|$
-$=\sum_{j=1}^{d}\sum_{k=1}^{d}|[C]_{kj}|\left(\frac{1}{n}\sum_{i=1}^{n}|[\bm{x}_{i}]_{k}|\right).$
+$\mathbb{E}_{\mathbb{P}_{n}}\left[\|C\bm{x}\|_{1}\right]=\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{d}|[C]_{j}\bm{x}_{i}|$ $\leq\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{d}\sum_{k=1}^{d}|[C]_{kj}||\bm{x}^{i}_{k}|$
+$=\sum_{j=1}^{d}\sum_{k=1}^{d}|[C]_{kj}|\left(\frac{1}{n}\sum_{i=1}^{n}|\bm{x}^{i}_{k}|\right).$
 We can conclude that the inequality holds. ∎
-A.4 Proof of Proposition 4.1
+Appendix D Proof for Theorem 3.4
+Our proof for Theorem 3.4 is based on the primal-dual witness (PDW) construction proposed by wainwright2009sharp; Ravikumar et al. (2011).
+Define the symmetric matrix $\hat{\Lambda}$ by $\hat{\Lambda}_{jj}=0$ and $\hat{\Lambda}_{kj}=\hat{\lambda}_{kj}$ for $k\neq j,k,j=1,\ldots,d$.
+The matrix $\widetilde{C}\succ 0$ is optimal for (10) if and only if there exists a symmetric $\widetilde{Z}$ such that
+$\bar{A}-\widetilde{C}^{-1}+\hat{\Lambda}\odot\widetilde{Z}=0,$ (17)
+where $\widetilde{Z}\in\partial\|\widetilde{C}\|_{1,1}$, $\bar{A}=\mathbb{E}_{\mathbb{P}_{n}}[\bm{x}\bm{x}^{\top}]$.
+Define the restricted problem
+$\dot{C}\in\arg\min_{C\succ 0,\ C_{S^{c}}=0}\Big\{-\log\det C+\langle\bar{A},C\rangle+\sum_{k\neq j}\hat{\lambda}_{kj}|C_{kj}|\Big\},$ (18)
+and set $\dot{Z}_{S}=\mathrm{sign}(\dot{C}_{S})$.
+For $(j,k)\in S^{c}$, we define
+$\dot{Z}_{jk}:=\frac{-\bar{A}_{jk}+{[{\dot{C}}^{-1}}]_{jk}}{\hat{\lambda}_{jk}}.$ (19)
+Let ${\it\Delta}:=\dot{C}-\Sigma^{-1}$, $W:=\bar{A}-\Sigma$, and
+$R({\it\Delta}):=\dot{C}^{-1}-\Sigma+\Sigma{\it\Delta}\Sigma.$
+We present several lemmas before further proceeding to prove Theorem 3.4.
+Lemma D.1 (Remainder Bound, lemma 5 in Ravikumar et al. (2011)).
+If
+$\|{\it\Delta}\|_{\max}\leq\frac{1}{3\kappa_{\Sigma}s},$
+then we have that
+$\|R({\it\Delta})\|_{\max}\leq\frac{3}{2}s\kappa_{\Sigma}^{3}\|{\it\Delta}\|_{\max}^{2}.$
+Lemma D.2 (Support error control).
+Define
+$\hat{\lambda}_{\max,S}:=\max_{(i,j)\in S}\hat{\lambda}_{ij},\qquad r:=2\kappa_{\Gamma}\Big(\|W\|_{\max}+\hat{\lambda}_{\max,S}\Big).$
+Assume that
+$r\leq\min\Big\{\frac{1}{3\kappa_{\Sigma}s},\ \frac{1}{3\kappa_{\Sigma}^{3}\kappa_{\Gamma}s}\Big\},$ (20)
+then the oracle error satisfies $\|\Delta\|_{\max}\leq r$.
+Proof.
+The restricted problem (18) is a convex program over a convex set.
+Moreover, the smooth part $C\mapsto-\log\det C+\langle S,C\rangle$ is strictly convex on $C\succ 0$.
+Therefore, the problem has a unique solution.
+We have ${\it\Delta}_{S^{c}}=0$ by PDW construction.
+Restricting the KKT condition of the restricted problem to $S$ and using $\bar{A}=\Sigma+W$ together with the remainder definition yields the following
+$\Big[W+\Sigma{\it\Delta}\Sigma-R({\it\Delta})+(\hat{\Lambda}\odot\dot{Z})\Big]_{S}=0.$ (21)
+We vectorize the left-hand side in (21) and use $\mathrm{vec}(\Sigma{\it\Delta}\Sigma)=\Gamma\mathrm{vec}({\it\Delta})$, and obtain the following
+$\Gamma_{SS}\mathrm{vec}({\it\Delta}_{S})-\mathrm{vec}\Big(R({\it\Delta})_{S}-W_{S}-(\hat{\Lambda}\odot\dot{Z})_{S}\Big)=0.$
+Multiplying the above by $(\Gamma_{SS})^{-1}$ gives
+$T({\it\Delta}):=\mathrm{vec}({\it\Delta}_{S})-(\Gamma_{SS})^{-1}\mathrm{vec}\Big(R({\it\Delta})_{S}-W_{S}-(\hat{\Lambda}\odot\dot{Z})_{S}\Big),$
+so that (21) holds if and only if $T({\it\Delta})=0$.
+For any $u\in\mathbb{R}^{|S|}$, let ${\it\Delta}(u)$ denote the symmetric matrix with ${\it\Delta}(u)_{S^{c}}=0$ and $\mathrm{vec}({\it\Delta}(u)_{S})=u$.
+Define $F:\mathbb{R}^{|S|}\to\mathbb{R}^{|S|}$ by
+$F(u):=u-T({\it\Delta}(u))=(\Gamma_{SS})^{-1}\mathrm{vec}\Big(R({\it\Delta}(u))_{S}-W_{S}-(\hat{\Lambda}\odot\dot{Z})_{S}\Big).$
+Then $F(u)=u$ if and only if $T({\it\Delta}(u))=0$, equivalently if and only if (21) holds for ${\it\Delta}(u)$.
+Let
+$\mathcal{B}(r):=\Big\{u\in\mathbb{R}^{|S|}:\ \|u\|_{\infty}\leq r\Big\}.$
+Take any $u\in\mathcal{B}(r)$ and , with a slight abuse of notation, write ${\it\Delta}={\it\Delta}(u)$.
+Because $\|\dot{Z}_{S}\|_{\max}\leq 1$ and the penalty is bounded by $\hat{\lambda}_{\max,S}$ on $S$, we have
+$\|(\hat{\Lambda}\odot\dot{Z})_{S}\|_{\max}\leq\hat{\lambda}_{\max,S}.$
+By definition of $\kappa_{\Gamma}=\|(\Gamma_{SS})^{-1}\|_{\infty}$, we obtain
+$\|F(u)\|_{\infty}\leq\kappa_{\Gamma}\Big(\|R({\it\Delta})\|_{\max}+\|W\|_{\max}+\hat{\lambda}_{\max,S}\Big).$
+Since $\|{\it\Delta}\|_{\max}\leq r$, $r\leq(3\kappa_{\Sigma}s)^{-1}$, and ${\it\Delta}_{S^{c}}=0$, Lemma D.1 indicates the remainder bound
+$\|R({\it\Delta})\|_{\max}\leq\frac{3}{2}s\kappa_{\Sigma}^{3}\|{\it\Delta}\|_{\max}^{2}\leq\frac{3}{2}s\kappa_{\Sigma}^{3}r^{2}.$
+Therefore,
+$\|F(u)\|_{\infty}\leq\kappa_{\Gamma}\Big(\tfrac{3}{2}s\kappa_{\Sigma}^{3}r^{2}+\|W\|_{\max}+\hat{\lambda}_{\max,S}\Big).$
+Using $r:=2\kappa_{\Gamma}(\|W\|_{\max}+\hat{\lambda}_{\max,S})$, we obtain
+$\|F(u)\|_{\infty}\leq\frac{3}{2}\kappa_{\Gamma}s\kappa_{\Sigma}^{3}r^{2}+\frac{r}{2}.$
+Finally, the condition $r\leq(3\kappa_{\Sigma}^{3}\kappa_{\Gamma}s)^{-1}$ implies
+$\frac{3}{2}\kappa_{\Gamma}s\kappa_{\Sigma}^{3}r^{2}\leq\frac{r}{2},$
+hence $\|F(u)\|_{\infty}\leq r$.
+Therefore, $F$ maps $\mathcal{B}(r)$ into itself.
+Since $F$ is continuous on $\mathcal{B}(r)$ and $\mathcal{B}(r)$ is convex and compact, the Brouwer’s fixed point theorem (ortega2000iterative) implies there exists $u^{\sharp}\in\mathcal{B}(r)$ such that $F(u^{\sharp})=u^{\sharp}$.
+Let ${\it\Delta}^{\sharp}={\it\Delta}(u^{\sharp})$, i.e.
+${\it\Delta}^{\sharp}_{S^{c}}=0$ and $\mathrm{vec}({\it\Delta}^{\sharp}_{S})=u^{\sharp}$.
+Then $F(u^{\sharp})=u^{\sharp}$ implies $T({\it\Delta}^{\sharp})=0$, so (21) holds for $\it\Delta$$^{\sharp}$.
+By the uniqueness of the restricted optimum, this fixed point must equal the solution ${\it\Delta}=\dot{C}-\Sigma^{-1}$.
+Therefore,
+$\|{\it\Delta}\|_{\max}=\|u^{\sharp}\|_{\infty}\leq r.$
+Lemma D.3 (Strict dual feasibility).
+Suppose Assumption 3.3 holds. Let $\dot{Z}$ be the PDW dual variable constructed in (18)–(19). Define
+$\hat{\lambda}_{\min,S^{c}}:=\min_{e\in S^{c}}\hat{\lambda}_{e},\quad\hat{\psi}:=\max_{e\in S^{c}}\sum_{s\in S}|A_{es}|\frac{\hat{\lambda}_{s}}{\hat{\lambda}_{e}}.$
+If $\hat{\psi}\leq 1-5\alpha/8$ and
+$\frac{1+\kappa_{A}}{\hat{\lambda}_{\min,S^{c}}}\Big(\|W\|_{\max}+\|R({\it\Delta})\|_{\max}\Big)\leq\frac{\alpha}{2},$ (22)
+then $\|\dot{Z}_{S^{c}}\|_{\max}<1$ and $\dot{C}=\widetilde{C}$.
+Proof.
+The vectorized KKT condition for the restricted problem (18) is
+$\Gamma{\it\Delta}+W-R({\it\Delta})+\hat{\Lambda}\dot{Z}=0.$ (23)
+We partition (23) into $(S,S^{c})$ blocks. Since our PDW construction enforces ${\it\Delta}_{S^{c}}=0$, we obtain
+$\Gamma_{SS}{\it\Delta}_{S}+W_{S}-R({\it\Delta})_{S}+\hat{\Lambda}_{S}\dot{Z}_{S}$ $=0,$ (24)
+$\Gamma_{S^{c}S}{\it\Delta}_{S}+W_{S^{c}}-R({\it\Delta})_{S^{c}}+\hat{\Lambda}_{S^{c}}\dot{Z}_{S^{c}}$ $=0.$ (25)
+Solving (24) gives
+${\it\Delta}_{S}=-(\Gamma_{SS})^{-1}\Big(W_{S}-R({\it\Delta})_{S}+\hat{\Lambda}_{S}\dot{Z}_{S}\Big).$ (26)
+Recall
+$A:=\Gamma_{S^{c}S}(\Gamma_{SS})^{-1}.$
+Substituting (26) into (25) yields
+$\hat{\Lambda}_{S^{c}}\dot{Z}_{S^{c}}=A\Big(W_{S}-R({\it\Delta})_{S}+\hat{\Lambda}_{S}\dot{Z}_{S}\Big)-\Big(W_{S^{c}}-R({\it\Delta})_{S^{c}}\Big).$ (27)
+Fix any $e\in S^{c}$. Taking absolute values of the $e$th component of (27) gives
+$|\dot{Z}_{e}|\leq\frac{1}{\hat{\lambda}_{e}}\Big(|[A(W_{S}-R({\it\Delta})_{S})]_{e}|+|W_{e}-R({\it\Delta})_{e}|+|[A\hat{\Lambda}_{S}\dot{Z}_{S}]_{e}|\Big).$ (28)
+Write row $e$ of $A$ as $(A_{es})_{s\in S}$. Then
+$|[A(W_{S}-R({\it\Delta})_{S})]_{e}|\leq\sum_{s\in S}|A_{es}|(|W_{s}|+|R({\it\Delta})_{s}|)\leq\|A_{e,\cdot}\|_{1}(\|W\|_{\max}+\|R({\it\Delta})\|_{\max}).$
+Also,
+$|W_{e}-R({\it\Delta})_{e}|\leq\|W\|_{\max}+\|R({\it\Delta})\|_{\max}.$
+Recall
+$\kappa_{A}=\max_{e\in S^{c}}\|A_{e,\cdot}\|_{1}.$
+Hence, uniformly over $e\in S^{c}$,
+$\frac{1}{\hat{\lambda}_{e}}\Big(|[A(W_{S}-R({\it\Delta})_{S})]_{e}|+|W_{e}-R({\it\Delta})_{e}|\Big)\leq\frac{1+\kappa_{A}}{\hat{\lambda}_{\min,S^{c}}}(\|W\|_{\max}+\|R({\it\Delta})\|_{\max}).$ (29)
+Since $|\dot{Z}_{s}|\leq 1$ for $s\in S$,
+$|[A\hat{\Lambda}_{S}\dot{Z}_{S}]_{e}|=\left|\sum_{s\in S}A_{es}\hat{\lambda}_{s}\dot{Z}_{s}\right|\leq\sum_{s\in S}|A_{es}|\hat{\lambda}_{s}.$
+Then
+$\frac{1}{\hat{\lambda}_{e}}|[A\hat{\Lambda}_{S}\dot{Z}_{S}]_{e}|\leq\hat{\psi}$. (30)
+Combining (29) and (30) in (28) and taking $\max_{e\in S^{c}}$ yields
+$\|\dot{Z}_{S^{c}}\|_{\max}\leq\frac{1+\kappa_{A}}{\hat{\lambda}_{\min,S^{c}}}\Big(\|W\|_{\max}+\|R({\it\Delta})\|_{\max}\Big)+\hat{\psi}$. (31)
+If (22) holds and $\hat{\psi}\leq 1-5\alpha/8$ hold, then
+$\|\dot{Z}_{S^{c}}\|_{\max}\leq\frac{\alpha}{2}+\left(1-\frac{5\alpha}{8}\right)<1.$
+Lemma D.4 (Sign recovery).
+If $\|\widetilde{C}-\Sigma^{-1}\|_{\max}\leq\min_{(i,j)\in E}|[\Sigma^{-1}]_{ij}|/2$, then $\mathrm{sign}(\widetilde{C}_{ij})=\mathrm{sign}([\Sigma^{-1}]_{ij})$ for all $(i,j)\in E$.
+Lemma D.5.
+If we have $\max_{1\leq i\leq d}\left|\hat{\omega}_{i}-\omega^{\ast}_{i}\right|\leq t_{\omega}$ and
+$\varepsilon_{n}:=\frac{2t_{\omega}}{\delta}<1,$
+then for all off-diagonal indices $s,e$,
+$\frac{\hat{\lambda}_{s}}{\hat{\lambda}_{e}}\leq\frac{1+\varepsilon_{n}}{1-\varepsilon_{n}}\frac{\lambda^{\ast}_{s}}{\lambda^{\ast}_{e}},\qquad\text{and hence}\qquad\hat{\psi}\leq\frac{1+\varepsilon_{n}}{1-\varepsilon_{n}}\psi^{\ast}.$
+In particular, if $\psi^{\ast}\leq 1-\alpha$ and $\varepsilon_{n}\leq\alpha/8$, then $\hat{\psi}\leq 1-5\alpha/8$.
+Proof.
+For any $i\neq j$,
+$|\hat{\lambda}_{ij}-\lambda^{\ast}_{ij}|=\delta\big|(\hat{\omega}_{i}-\omega^{\ast}_{i})+(\hat{\omega}_{j}-\omega^{\ast}_{j})\big|\leq 2\delta t_{\omega}.$
+Also $\lambda^{\ast}_{ij}\geq\delta^{2}$, we have
+$\frac{|\hat{\lambda}_{ij}-\lambda^{\ast}_{ij}|}{\lambda^{\ast}_{ij}}\leq\frac{2\delta t_{\omega}}{\delta^{2}}=\frac{2t_{\omega}}{\delta}=\varepsilon_{n}.$
+Thus,
+$(1-\varepsilon_{n})\lambda^{\ast}_{ij}\leq\hat{\lambda}_{ij}\leq(1+\varepsilon_{n})\lambda^{\ast}_{ij},$
+implying
+$\frac{\hat{\lambda}_{s}}{\hat{\lambda}_{e}}\leq\frac{1+\varepsilon_{n}}{1-\varepsilon_{n}}\frac{\lambda_{s}^{\ast}}{\lambda_{e}^{\ast}}.$
+Taking maxima in the definition of $\hat{\psi},\psi^{\ast}$ yields the inequality
+$\hat{\psi}\leq\frac{1+\varepsilon_{n}}{1-\varepsilon_{n}}\psi^{\ast}.$
+If $\varepsilon_{n}\leq\alpha/8$, then
+$\frac{1+\varepsilon_{n}}{1-\varepsilon_{n}}\leq 1+3\varepsilon_{n}\leq 1+\frac{3\alpha}{8},$
+and therefore
+$\hat{\psi}\leq\left(1+\frac{3\alpha}{8}\right)(1-\alpha)\leq 1-\frac{5\alpha}{8}.$
+Lemma D.6 (Concentration Inequality I).
+There exist universal constants $C_{1}>0,c_{1}>0$ such that for any $\tau>2$,
+$\mathbb{P}\left(\|W\|_{\max}\leq C_{1}\sigma_{\max}^{2}\sqrt{\frac{\tau\log d}{n}}\right)\ \geq\ 1-2d^{2-\tau}$
+holds when $n\geq c_{1}\tau\log d.$
+Proof.
+Let $Y_{t}:=\bm{x}^{t}_{i}\bm{x}^{t}_{j}-\Sigma_{ij}$.
+Under Gaussianity, $Y_{t}$ is sub-exponential with $\|Y_{t}\|_{\psi_{1}}\leq K_{1}\sigma_{\max}^{2}$ for a universal $K_{1}$.
+The Bernstein concentration inequality for sub-exponential variables gives
+$\mathbb{P}\left(\left|\frac{1}{n}\sum_{t=1}^{n}Y_{t}\right|\geq u\right)\leq 2\exp\!\left(-C_{0}n\min\Big\{\frac{u^{2}}{K_{1}^{2}\sigma_{\max}^{4}},\frac{u}{K_{1}\sigma_{\max}^{2}}\Big\}\right).$ (32)
+We take
+$u=C_{1}\sigma_{\max}^{2}\sqrt{\frac{\tau\log d}{n}}.$
+If
+$n\geq\Big(\frac{C_{1}}{K_{1}}\Big)^{2}\tau\log d,$
+we have $u\leq K_{1}\sigma_{\max}^{2}$, so the minimum in the concentration inequality (32) is achieved by the quadratic term. Hence
+$\mathbb{P}(|W_{ij}|\geq u)\leq 2\exp\!\left(-C_{0}n\frac{u^{2}}{K_{1}^{2}\sigma_{\max}^{4}}\right)=2\exp\!\left(-C_{0}\frac{C_{1}^{2}}{K_{1}^{2}}\tau\log d\right).$
+Choosing $C_{1}\geq K_{1}/\sqrt{C_{0}}$ results in $\mathbb{P}(|W_{ij}|\geq u)\leq 2d^{-\tau}$. The union bound over $d^{2}$ pairs gives the result.
+Lemma D.7 (Concentration Inequality II).
+There exist a universal constant $C_{2}>0$ such that for any $\tau>1$,
+$\mathbb{P}\!\left(\max_{1\leq i\leq d}\left|\mathbb{E}_{\mathbb{P}_{n}}[|\bm{x}_{i}|]-\mathbb{E}[|\bm{x}_{i}|]\right|\leq C_{2}\sigma_{\max}\sqrt{\frac{\tau\log d}{n}}\right)\ \geq\ 1-2d^{1-\tau}.$
+Proof.
+Let $Z_{t}:=|\bm{x}^{t}_{i}|-\mathbb{E}[|\bm{x}_{i}|]$. $Z_{t}$ is a sub-Gaussian random variable with $\|Z_{t}\|_{\psi_{2}}\leq K_{2}\sigma_{\max}$ for some universal constant $K_{2}$.
+We use the Hoeffding-type concentration bound:
+$\mathbb{P}\left(\left|\frac{1}{n}\sum_{t=1}^{n}Z_{t}\right|\geq u\right)\leq 2\exp\!\left(-\frac{cnu^{2}}{K_{2}^{2}\sigma_{\max}^{2}}\right).$
+Take
+$u=C_{2}\sigma_{\max}\sqrt{\frac{\tau\log d}{n}},\quad C_{2}=\frac{K_{2}}{\sqrt{c}},$
+and substitute into the exponent:
+$\frac{cn}{K_{2}^{2}\sigma_{\max}^{2}}\left(C_{2}^{2}\sigma_{\max}^{2}\frac{\tau\log d}{n}\right)=-\tau\log d.$
+Applying the union bound over $d$ indices gives the result. ∎
+Equipped with the lemmas above, we proceed to prove Theorem 3.4.
+Proof.
+Define the event
+$\mathcal{E}:=\left\{\|W\|_{\max}\leq C_{1}\sigma_{\max}^{2}\sqrt{\frac{\tau\log d}{n}}\right\}\cap\left\{\max_{i}|\hat{\omega}_{i}-\omega_{i}^{\ast}|\leq C_{2}\sigma_{\max}\sqrt{\frac{\tau\log d}{n}}\right\}.$
+By Lemma D.6, Lemma D.7 and a union bound,
+$\mathbb{P}(\mathcal{E})\geq 1-2d^{2-\tau}-2d^{1-\tau}\geq 1-4d^{2-\tau}.$
+We complete our proof in six steps stated below.
+Step 1: control $\hat{\psi}$.
+On the event $\mathcal{E}$, we define
+$t_{\omega}:=C_{2}\sigma_{\max}\sqrt{\frac{\tau\log d}{n}},\qquad\varepsilon_{n}:=\frac{2t_{\omega}}{\delta}=\frac{2C_{2}\sigma_{\max}\sqrt{\tau}}{c_{\delta}}.$
+By the definition of $c_{\delta}$, we have
+$\varepsilon_{n}\leq\frac{\alpha}{8},$
+then Lemma D.5 implies
+$\hat{\psi}\leq 1-\frac{5\alpha}{8}.$
+Step 2: lower bound $\hat{\lambda}_{\min,S^{c}}$.
+For any off-diagonal $e=(i,j)\in S^{c}$,
+$\hat{\lambda}_{e}=\delta(\hat{\omega}_{i}+\hat{\omega}_{j})+\delta^{2}\geq\delta\big((\omega_{i}^{\ast}-t_{\omega})+(\omega_{j}^{\ast}-t_{\omega})\big)+\delta^{2}\geq 2\delta(\omega_{\min}^{\ast}-t_{\omega})+\delta^{2}.$
+Since we have
+$n\geq 2\pi C_{2}^{2}\left(\frac{\sigma_{\max}}{\sigma_{\min}}\right)^{2}\tau\log d,$
+then
+$t_{\omega}=C_{2}\sigma_{\max}\sqrt{\frac{\tau\log d}{n}}\leq\frac{1}{2}\sqrt{\frac{2}{\pi}}\sigma_{\min}=\frac{\omega_{\min}^{\ast}}{2},$
+and therefore
+$\hat{\lambda}_{\min,S^{c}}:=\min_{e\in S^{c}}\hat{\lambda}_{e}\geq\ \omega_{\min}^{\ast}\delta.$ (33)
+Step 3: upper bound $\hat{\lambda}_{\max,S}$.
+$\hat{\lambda}_{ij}=\delta(\hat{\omega}_{i}+\hat{\omega}_{j})+\delta^{2}\leq 2\delta(\omega_{\max}^{\ast}+t_{\omega})+\delta^{2}.$
+We have $t_{\omega}\leq\omega_{\min}^{\ast}/2\leq\omega_{\max}^{\ast}/2$, hence
+$\hat{\lambda}_{\max,S}:=\max_{(i,j)\in S}\hat{\lambda}_{ij}\leq 3\omega_{\max}^{\ast}\delta+\delta^{2}.$
+Combining with $\|W\|_{\max}\leq C_{1}\sigma_{\max}^{2}\sqrt{\tau\log d/n}$, $\delta=c_{\delta}\sqrt{\log d/n}$, and $n\geq\log d$, we obtain
+$\|W\|_{\max}+\hat{\lambda}_{\max,S}\leq\Big(C_{1}\sigma_{\max}^{2}\sqrt{\tau}+3\omega_{\max}^{\ast}c_{\delta}+c_{\delta}^{2}\Big)\sqrt{\frac{\log d}{n}}=B\sqrt{\frac{\log d}{n}}.$
+Define as in Lemma D.2
+$r:=2\kappa_{\Gamma}\Big(\|W\|_{\max}+\hat{\lambda}_{\max,S}\Big),$
+then we can have that
+[…]
+Step 4: support error and remainder control.
+The sample size condition and (34) implies
+$r\leq\frac{1}{3\kappa_{\Sigma}s},\quad r\leq\frac{1}{3\kappa_{\Sigma}^{3}\kappa_{\Gamma}s},$
+so the condition (20) in Lemma D.2 holds.
+Hence, the oracle error follows
+$\|{\it\Delta}\|_{\max}\leq r\leq 2\kappa_{\Gamma}B\sqrt{\frac{\log d}{n}}.$ (35)
+Then, Lemma D.1 gives
+$\|R({\it\Delta})\|_{\max}\leq\frac{3}{2}s\kappa_{\Sigma}^{3}\|\Delta\|_{\max}^{2}\leq\frac{3}{2}s\kappa_{\Sigma}^{3}r^{2}\leq 6s\kappa_{\Sigma}^{3}\kappa_{\Gamma}^{2}B^{2}\frac{\log d}{n}.$ (36)
+Step 5: strict dual feasibility.
+By Lemma D.3, it suffices to check:
+$\hat{\psi}\leq 1-\frac{5\alpha}{8},\quad\frac{1+\kappa_{A}}{\hat{\lambda}_{\min,S^{c}}}\Big(\|W\|_{\max}+\|R({\it\Delta})\|_{\max}\Big)\leq\frac{\alpha}{2}.$
+The first inequality is proved in Step 1.
+For the second inequality, we split the left-hand side into the $W$-term and the $R$-term.
+(a) It follows from (33) and $\|W\|_{\max}\leq C_{1}\sigma_{\max}^{2}\sqrt{\tau\log d/n}$ that
+$\frac{1+\kappa_{A}}{\hat{\lambda}_{\min,S^{c}}}\|W\|_{\max}\leq\frac{1+\kappa_{A}}{\omega_{\min}^{\ast}\delta}C_{1}\sigma_{\max}^{2}\sqrt{\frac{\tau\log d}{n}}=\frac{(1+\kappa_{A})C_{1}\sigma_{\max}^{2}\sqrt{\tau}}{\omega_{\min}^{\ast}c_{\delta}}.$
+Since $\omega_{\min}^{\ast}=\sqrt{2/\pi}\sigma_{\min}$, and by the definition of $c_{\delta}$, we get
+$\frac{(1+\kappa_{A})C_{1}\sigma_{\max}^{2}\sqrt{\tau}}{\omega_{\min}^{\ast}c_{\delta}}\leq\frac{\alpha}{4}.$
+(b) It follows from (33) and (36) that
+$\frac{1+\kappa_{A}}{\hat{\lambda}_{\min,S^{c}}}\|R({\it\Delta})\|_{\max}\leq\frac{6(1+\kappa_{A})s\kappa_{\Sigma}^{3}\kappa_{\Gamma}^{2}B^{2}}{\omega_{\min}^{\ast}c_{\delta}}\sqrt{\frac{\log d}{n}}.$
+The condition of the sample size $n$ implies
+$\sqrt{\frac{\log d}{n}}\leq\frac{\alpha\omega_{\min}^{\ast}c_{\delta}}{24(1+\kappa_{A})s\kappa_{\Sigma}^{3}\kappa_{\Gamma}^{2}B^{2}},$
+and
+$\frac{6(1+\kappa_{A})s\kappa_{\Sigma}^{3}\kappa_{\Gamma}^{2}B^{2}}{\omega_{\min}^{\ast}c_{\delta}}\sqrt{\frac{\log d}{n}}\leq\frac{\alpha}{4}.$
+Combining (a) and (b) indicates
+$\frac{1+\kappa_{A}}{\hat{\lambda}_{\min,S^{c}}}\Big(\|W\|_{\max}+\|R({\it\Delta})\|_{\max}\Big)\leq\frac{\alpha}{2}.$
+Therefore, we apply Lemma D.3 and obtain strict dual feasibility $\|\dot{Z}_{S^{c}}\|_{\max}<1$, implying the PDW solution coincides with the global optimum:
+$\widetilde{C}=\dot{C}\quad\text{and}\quad\widetilde{C}_{S^{c}}=0.$
+Step 6: convergence rate and sign recovery.
+Since $\widetilde{C}=\dot{C}$ and $\dot{C}-\Sigma^{-1}={\it\Delta}$, it follows from (35) that,
+$\|\widetilde{C}-\Sigma^{-1}\|_{\max}=\|{\it\Delta}\|_{\max}\leq 2\kappa_{\Gamma}B\sqrt{\frac{\log d}{n}}.$
+Finally, if
+$\min_{(i,j)\in E}|[\Sigma^{-1}]_{ij}|\geq 4\kappa_{\Gamma}B\sqrt{\frac{\log d}{n}},$
+then $\|\widetilde{C}-\Sigma^{-1}\|_{\max}\leq\min_{(i,j)\in E}|[\Sigma^{-1}]_{ij}|/2$, and Lemma D.4 yields sign recovery. ∎
+Appendix E Proof of Proposition 3.5
+Proof.
+By the property of the Kronecker product,
+$\Gamma=\Sigma\otimes\Sigma=(DRD)\otimes(DRD)=(D\otimes D)(R\otimes R)(D\otimes D)=(D\otimes D)\Gamma^{\dagger}(D\otimes D).$
+Define $D_{\Gamma}:=D\otimes D$. Then, we have that
+$\Gamma_{S^{c}S}=(D_{\Gamma})_{S^{c}S^{c}}\Gamma^{\dagger}_{S^{c}S}(D_{\Gamma})_{SS},\quad\Gamma_{SS}=(D_{\Gamma})_{SS}\Gamma^{\dagger}_{SS}(D_{\Gamma})_{SS}.$
+Since $(D_{\Gamma})_{SS}$ is diagonal and invertible, we have that
+$(\Gamma_{SS})^{-1}=\big((D_{\Gamma})_{SS}\Gamma^{\dagger}_{SS}(D_{\Gamma})_{SS}\big)^{-1}=(D_{\Gamma})_{SS}^{-1}(\Gamma^{\dagger}_{SS})^{-1}(D_{\Gamma})_{SS}^{-1},$
+$A=\Gamma_{S^{c}S}(\Gamma_{SS})^{-1}=(D_{\Gamma})_{S^{c}S^{c}}\Gamma^{\dagger}_{S^{c}S}(\Gamma^{\dagger}_{SS})^{-1}(D_{\Gamma})_{SS}^{-1}=(D_{\Gamma})_{S^{c}S^{c}}A^{\dagger}(D_{\Gamma})_{SS}^{-1}.$ (37)
+Let $d_{u}:=\big[D_{\Gamma}\big]_{uu}$. It follows from (37) that
+$A_{es}=d_{e}A^{\dagger}_{es}d_{s}^{-1}.$
+Hence,
+$|A_{es}|=|A^{\dagger}_{es}|\frac{d_{e}}{d_{s}}.$
+Therefore,
+$\max_{e\in S^{c}}\sum_{s\in S}|A_{es}|\leq\max_{e\in S^{c}}d_{e}\max_{s\in S}d_{s}^{-1}\|A^{\dagger}\|_{\infty}.$
+Since we have $d_{u}=\sqrt{\Sigma_{ii}}\sqrt{\Sigma_{jj}}$, we have that
+$\max_{e\in S^{c}}\sum_{s\in S}|A_{es}|\leq\Big(\frac{\sigma_{\max}}{\sigma_{\min}}\Big)^{2}\|A^{\dagger}\|_{\infty},$
+which proves the first inequality.
+By definition, we have that
+$\psi^{\ast}=\max_{e\in S^{c}}\sum_{s\in S}|A_{es}|\frac{\lambda_{s}^{\ast}}{\lambda_{e}^{\ast}}=\max_{e\in S^{c}}\sum_{s\in S}|A^{\dagger}_{es}|\frac{d_{e}}{d_{s}}\frac{\lambda_{s}^{\ast}}{\lambda_{e}^{\ast}}.$
+If we define $w_{u}:=\lambda_{u}^{\ast}/d_{u}$, then we have that
+$\psi^{\ast}=\max_{e\in S^{c}}\sum_{s\in S}|A^{\dagger}_{es}|\frac{w_{s}}{w_{e}}\leq\frac{\max_{s\in S}w_{s}}{\min_{e\in S^{c}}w_{e}}\max_{e\in S^{c}}\sum_{s\in S}|A^{\dagger}_{es}|=\frac{\max_{s\in S}w_{s}}{\min_{e\in S^{c}}w_{e}}\|A^{\dagger}\|_{\infty}.$ (38)
+Recall that for $i\neq j$,
+$\lambda_{ij}^{\ast}=\delta(\omega_{i}^{\ast}+\omega_{j}^{\ast})+\delta^{2},\qquad\omega_{i}^{\ast}=\sqrt{2/\pi}\sqrt{\Sigma_{ii}}.$
+Then for $i\neq j$,
+$w_{(i,j)}=\frac{\lambda_{ij}^{\ast}}{\sqrt{\Sigma_{ii}}\sqrt{\Sigma_{jj}}}=\delta\sqrt{\frac{2}{\pi}}\Big(\frac{1}{\sqrt{\Sigma_{ii}}}+\frac{1}{\sqrt{\Sigma_{jj}}}\Big)+\frac{\delta^{2}}{\sqrt{\Sigma_{ii}}\sqrt{\Sigma_{jj}}}.$
+Using $\sigma_{\min}\leq\sqrt{\Sigma_{ii}},\sqrt{\Sigma_{jj}}\leq\sigma_{\max}$, we have the bounds
+$w_{(i,j)}\leq\delta\sqrt{\frac{2}{\pi}}\frac{2}{\sigma_{\min}}+\frac{\delta^{2}}{\sigma_{\min}^{2}},\qquad w_{(i,j)}\geq\delta\sqrt{\frac{2}{\pi}}\frac{2}{\sigma_{\max}}+\frac{\delta^{2}}{\sigma_{\max}^{2}}.$
+Therefore,
+$\frac{\max_{s\in S}w_{s}}{\min_{e\in S^{c}}w_{e}}\leq\Big(\frac{\sigma_{\max}}{\sigma_{\min}}\Big)^{2}\frac{2\delta\sqrt{\frac{2}{\pi}}\sigma_{\min}+\delta^{2}}{2\delta\sqrt{\frac{2}{\pi}}\sigma_{\max}+\delta^{2}}.$
+Since $10\delta<\sqrt{2/\pi}\sigma_{\min}$,
+$2\delta\sqrt{\frac{2}{\pi}}\sigma_{\min}+\delta^{2}\leq 21\delta\sqrt{\frac{2}{\pi}}\sigma_{\min}/10,\qquad 2\delta\sqrt{\frac{2}{\pi}}\sigma_{\max}+\delta^{2}\geq 2\delta\sqrt{\frac{2}{\pi}}\sigma_{\max},$
+and hence
+$\frac{2\delta\sqrt{\frac{2}{\pi}}\sigma_{\min}+\delta^{2}}{2\delta\sqrt{\frac{2}{\pi}}\sigma_{\max}+\delta^{2}}\leq\frac{21}{20}\frac{\sigma_{\min}}{\sigma_{\max}}.$
+Then, we have that
+$\frac{\max_{s\in S}w_{s}}{\min_{e\in S^{c}}w_{e}}\leq\frac{21}{20}\frac{\sigma_{\max}}{\sigma_{\min}}.$
+Plugging into (38) gives
+$\psi^{\ast}\leq\frac{21}{20}\Big(\frac{\sigma_{\max}}{\sigma_{\min}}\Big)\|A^{\dagger}\|_{\infty},$
+which proves the second inequality. ∎
+Appendix F Proof of Proposition 4.1
 We first state two propositions before we prove Proposition 4.1.
-Proposition A.2.
+Proposition F.1.
 For vector $b\in\mathbb{R}^{d}$, positive definite matrix $Q\in\mathbb{R}^{d\times d}$ and $p\in(1,\infty)$, we have that
@@
 Since the constraint set $\|\Delta\|_{p}\leq 1$ is convex, it follows from the KKT condition that the solution to the problem, denoted by $\Delta_{\ast}$ and $\lambda_{\ast}$, follows the equations:
-$\|\Delta_{\ast}\|_{p}^{p}-1=0,\quad b+\varepsilon 2Q\Delta_{\ast}-\lambda_{\ast}p\|\Delta_{\ast}\|_{p}^{p-1}|\Delta_{\ast}|^{p-1}\odot\operatorname{sign}(\Delta_{\ast})=0.$ (18)
+$\|\Delta_{\ast}\|_{p}^{p}-1=0,\quad b+\varepsilon 2Q\Delta_{\ast}-\lambda_{\ast}p\|\Delta_{\ast}\|_{p}^{p-1}|\Delta_{\ast}|^{p-1}\odot\mathrm{sign}(\Delta_{\ast})=0.$ (40)
 Assume there are two distinct solutions to problem (39) when $\varepsilon\to 0$, denoted by $\Delta_{\ast}^{1}$ and $\Delta_{\ast}^{2}$.
 Equivalently, both $\Delta_{\ast}^{1}$ and $\Delta_{\ast}^{2}$ satisfy the equations in (40), indicating
-$2\varepsilon Q(\Delta_{\ast}^{1}-\Delta_{\ast}^{2})=\lambda_{\ast}p(|\Delta_{\ast}^{2}|^{p-1}\odot\operatorname{sign}(\Delta_{\ast}^{2})-|\Delta_{\ast}^{1}|^{p-1}\odot\operatorname{sign}(\Delta_{\ast}^{1})),$ (19)
+$2\varepsilon Q(\Delta_{\ast}^{1}-\Delta_{\ast}^{2})=\lambda_{\ast}p(|\Delta_{\ast}^{2}|^{p-1}\odot\mathrm{sign}(\Delta_{\ast}^{2})-|\Delta_{\ast}^{1}|^{p-1}\odot\mathrm{sign}(\Delta_{\ast}^{1})),$ (41)
 $\|\Delta_{\ast}^{1}\|_{p}=\|\Delta_{\ast}^{2}\|_{p}=1,\quad\Delta_{\ast}^{1}\not=\Delta_{\ast}^{2}$
@@
 as $\varepsilon\to 0$. ∎
-Proposition A.3.
-For vector $b\in\mathbb{R}^{d}$ with $b]_{i}\not=0$ for i=1,,di=1,...,d, positive definite matrix $Q\in\mathbb{R}^{d\times d}$, we have that
-$\max_{\|\Delta\|_{\infty}\leq 1}b^{\top}\Delta+\varepsilon\Delta^{\top}Q\Delta=\|b\|_{1}+\varepsilon\operatorname{sign}(b)^{\top}Q\operatorname{sign}(b),$ (20)
+Proposition F.2.
+For vector $b\in\mathbb{R}^{d}$ with $b]_{i}\not=0$ for $i=1,\ldots,d$, positive definite matrix $Q\in\mathbb{R}^{d\times d}$, we have that
+$\max_{\|\Delta\|_{\infty}\leq 1}b^{\top}\Delta+\varepsilon\Delta^{\top}Q\Delta=\|b\|_{1}+\varepsilon\mathrm{sign}(b)^{\top}Q\mathrm{sign}(b),$ (42)
 as $\varepsilon\to 0$.
 Proof.
-When $\varepsilon=0$, the optimization problem $\max_{\|\Delta\|_{\infty}\leq 1}b^{\top}\Delta$ has the unique solution at $\operatorname{sign}(b)$.
+When $\varepsilon=0$, the optimization problem $\max_{\|\Delta\|_{\infty}\leq 1}b^{\top}\Delta$ has the unique solution at $\mathrm{sign}(b)$.
 When $\varepsilon\not=0$, the objective function is convex and the constraint set is a convex compact set.
@@
 Since $b]_{i}\not=0$, when $\varepsilon$ is very small, flipping the sign of $\Delta^{\ast}$ will decrease the value of the objective function.
-Thus, the optimal solution is $\operatorname{sign}(b)$ when $\varepsilon\to 0$.
+Thus, the optimal solution is $\mathrm{sign}(b)$ when $\varepsilon\to 0$.
 ∎
@@
 $\max_{\|\bm{\Delta}\|_{p}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C(\bm{x}+\bm{\Delta})=\bm{x}^{\top}C\bm{x}+\delta\max_{\|\bm{\Delta}\|_{p}\leq 1}\left\{2(C\bm{x})^{\top}\bm{\Delta}+\delta\bm{\Delta}^{\top}C\bm{\Delta}\right\}.$
-In the probability space $(\bm{P},C,\mathcal{F})$ for variable $\bm{x}$, for each $\bm{\omega}\in C$, it follows from Proposition A.2 and Proposition A.3 that
+In the probability space $(\bm{P},C,\mathcal{F})$ for variable $\bm{x}$, for each $\bm{\omega}\in C$, it follows from Proposition F.1 and Proposition F.2 that
 $\max_{\|\bm{\Delta}\|_{p}\leq 1}\left\{2(C\bm{x}(\bm{\omega}))^{\top}\bm{\Delta}+\delta\bm{\Delta}^{\top}C\bm{\Delta}\right\}=2\|C\bm{x}(\bm{\omega})\|_{q}+\delta\bm{v}_{\bm{x}(\bm{\omega})}^{\top}C\bm{v}_{\bm{x}(\bm{\omega})}+\mathcal{R}(\bm{\omega}),$
@@
 $\lim_{\delta\to 0}\frac{\mathbb{E}_{\bm{P}}[\mathcal{R}]}{\delta}=\int_{\bm{\omega}\in C}\lim_{\delta\to 0}\frac{\mathcal{R}(\bm{\omega})}{\delta}d\bm{P}(\bm{\omega})=0.$
-A.5 Proof of Theorem 5.1
+Appendix G Proof of Theorem 5.1
 Proof.
@@
 where $\mathrm{vec}(\mathbf{G})\sim\mathcal{N}(\bm{0},\Lambda)$, and $\Lambda$ is such that
-$\mathrm{cov}([\mathbf{G}]_{ij},[\mathbf{G}]_{i^{\prime}j^{\prime}})=\mathrm{cov}([\bm{x}]_{i}[\bm{x}]_{j},[\bm{x}]_{i^{\prime}}[\bm{x}]_{j^{\prime}}).$
+$\mathrm{cov}([\mathbf{G}]_{ij},[\mathbf{G}]_{i^{\prime}j^{\prime}})=\mathrm{cov}(\bm{x}_{i}\bm{x}_{j},\bm{x}_{i^{\prime}}\bm{x}_{j^{\prime}}).$
 Notice that we also have
@@
 $\sqrt{n}\left(\widehat{C}-\Sigma^{-1}\right)\Rightarrow-2\eta\Sigma^{-1}\mathbb{E}\left[\frac{\left(\mathrm{sign}(\Sigma^{-1}\bm{x})\odot|\Sigma^{-1}\bm{x}|^{q-1}\right)\bm{x}^{\top}}{\|\Sigma^{-1}\bm{x}\|_{q}^{q-1}}\right]\Sigma^{-1}.$
-A.6 Proof of Theorem 5.2
+Appendix H Proof of Theorem 5.2
 Proof.
 We define
-$\Phi_{n}(C)=-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{x}^{\top}C\bm{x}\right]+2\delta\sum_{j=1}^{d}\sum_{k=1}^{d}\left(\frac{1}{n}\sum_{i=1}^{n}|[\bm{x}_{i}]_{k}|\right)|[C]_{kj}|+\delta^{2}\|C\|_{1,1}.$
+$\Phi_{n}(C)=-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{x}^{\top}C\bm{x}\right]+2\delta\sum_{j=1}^{d}\sum_{k=1}^{d}\left(\frac{1}{n}\sum_{i=1}^{n}|\bm{x}^{i}_{k}|\right)|[C]_{kj}|+\delta^{2}\|C\|_{1,1}.$
 For $C\succ 0$, the function $\Phi_{n}(C)$ is convex.
@@
 $=$ $-n\log\det(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U})+n\log\det(\Sigma^{-1})+\sqrt{n}\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{x}^{\top}\bm{U}\bm{x}\right]$
-$+2\eta\sqrt{n}\sum_{j=1}^{d}\sum_{k=1}^{d}\left(\frac{1}{n}\sum_{i=1}^{n}|[\bm{x}_{i}]_{k}|\right)\left(|[\Sigma^{-1}]_{kj}+\frac{1}{\sqrt{n}}[\bm{U}]_{kj}|-|[\Sigma^{-1}]_{kj}|\right)$
+$+2\eta\sqrt{n}\sum_{j=1}^{d}\sum_{k=1}^{d}\left(\frac{1}{n}\sum_{i=1}^{n}|\bm{x}^{i}_{k}|\right)\left(|[\Sigma^{-1}]_{kj}+\frac{1}{\sqrt{n}}\bm{U}_{kj}|-|[\Sigma^{-1}]_{kj}|\right)$
 $+o(1).$
@@
 $W(\bm{U})$ $=-\mathrm{tr}(\bm{U}\mathbf{G})+\frac{1}{2}\mathrm{tr}(\bm{U}\Sigma\bm{U}\Sigma)$
-$+2\eta\sum_{j=1}^{d}\sum_{k=1}^{d}\mathbb{E}[|[\bm{x}]_{k}|]\left([\bm{U}]_{jk}\operatorname{sign}([\Sigma^{-1}]_{jk})\mathbb{I}([\Sigma^{-1}]_{jk}\not=0)+|[\bm{U}]_{jk}|\mathbb{I}([\Sigma^{-1}]_{jk}=0)\right).$
+$+2\eta\sum_{j=1}^{d}\sum_{k=1}^{d}\mathbb{E}[|\bm{x}_{k}|]\left(\bm{U}_{jk}\mathrm{sign}([\Sigma^{-1}]_{jk})\mathbb{I}([\Sigma^{-1}]_{jk}\not=0)+|\bm{U}_{jk}|\mathbb{I}([\Sigma^{-1}]_{jk}=0)\right).$
 Case 2: $\gamma>1/2$.
@@
 $W(\bm{U})$ $=\frac{1}{2}\mathrm{tr}(\bm{U}\Sigma\bm{U}\Sigma)$
-$+2\eta\sum_{j=1}^{d}\sum_{k=1}^{d}\mathbb{E}[|[\bm{x}]_{k}|]\left([\bm{U}]_{jk}\operatorname{sign}([\Sigma^{-1}]_{jk})\mathbb{I}([\Sigma^{-1}]_{jk}\not=0)+|[\bm{U}]_{jk}|\mathbb{I}([\Sigma^{-1}]_{jk}=0)\right).$
-Since we have $\bm{x}\sim\mathcal{N}(\bm{0},\Sigma)$, then we have $\bm{x}]_{k}\sim\mathcal{N}(\bm{0},[\Sigma]_{kk})$. In this way,
-$\mathbb{E}[|[\bm{x}]_{k}|]=\sqrt{\frac{2[\Sigma]_{kk}}{\pi}}.$
+$+2\eta\sum_{j=1}^{d}\sum_{k=1}^{d}\mathbb{E}[|\bm{x}_{k}|]\left(\bm{U}_{jk}\mathrm{sign}([\Sigma^{-1}]_{jk})\mathbb{I}([\Sigma^{-1}]_{jk}\not=0)+|\bm{U}_{jk}|\mathbb{I}([\Sigma^{-1}]_{jk}=0)\right).$
+Since we have $\bm{x}\sim\mathcal{N}(\bm{0},\Sigma)$, then we have $\bm{x}_{k}\sim\mathcal{N}(\bm{0},\Sigma_{kk})$. In this way,
+$\mathbb{E}[|\bm{x}_{k}|]=\sqrt{\frac{2\Sigma_{kk}}{\pi}}.$